### v1

# Scaling Up Active Testing to Large Language Models

Abstract
Active testing enables label-efficient evaluation of models through careful data acquisition. However, its significant computational costs have previously undermined its use for large models. We show how it can be successfully scaled up to the evaluation of large language models (LLMs). In particular we show that the surrogate model used to guide data acquisition can be constructed cheaply using in-context learning, does not require updating within an active-testing loop, and can be smaller than the target model. We even find we can make good data-acquisition decisions without computing predictions with the target model and further introduce a single-run error estimator to asses how well active testing is working on the fly. We find that our approach is able to more effectively evaluate LLM performance with less data than current standard practices.
1 Introduction
Evaluating frontier models is becoming more expensive as they become more sophisticated [5, 25] and currently represents a critical challenge in the field [27, 36, 39]. Moreover, with new models arriving in quick succession, and ever-present scope for data to leak from evaluations to training [9], the evaluation problem is a dynamic one: it requires ongoing, adaptive gathering of new evaluation data.
Active testing [16, 17] is an attractive solution to this problem. Motivated by the observation that some labels are more informative than others about the behaviour of a target model, it involves carefully, and typically adaptively, deciding which test inputs to acquire labels for. The foundation for this is a surrogate model of the test-time label distribution, which is typically updated on the labels acquired during testing. Making a data-acquisition decision requires computing predictions with the surrogate model and often also the target model. The computational cost of this process is often justifiable in light of high labelling costs—but it limits the scope of active testing’s applicability.
In this work, we show how active testing can be scaled up to LLMs, thereby allowing them to be more efficiently evaluated with less labelled test data. To this end, we first highlight three potential computational bottlenecks in the active testing pipeline: updating the surrogate model as labels are acquired, computing predictions with the surrogate model, and computing predictions with the target model. The first of these is particularly prominent: it has historically involved repeatedly running gradient-based training on the acquired test data within an active-testing loop.
Figure 1: Our proposed active-testing approach enables low-error estimates of the risk (expected predictive test loss) of a large language model (here the 70B version of Llama 2) on four text-classification problems (SST-2, FPB, HS and Subj) while only using a small label budget. We compare uniform-random testing against active testing (LURE), using either a 7B or 70B surrogate model with in-context learning to guide data acquisition.
We then identify simple and surprisingly effective ways to overcome each of these key bottlenecks. First, we strip back the iterative training of the underlying surrogate-model to a single step of in-context learning [4] on a small amount of initial test data, removing the need for repeated gradient-based training. Second, we show it is possible to use a surrogate model that is smaller than the target model, reducing the cost of surrogate-model predictions. Third, we show that one can even forgo computing predictions with the target model and rely solely on the surrogate model for data acquisition. Together these changes allow us to scale up active testing to LLMs.
Empirically, our approach substantially improves over the standard practice of acquiring data at random (Figure˜1), typically reducing estimation error by 25% to 50%, and sometimes up to 80%.
To further improve the real-world applicability of active testing, we also address the challenge of judging how well it is working on a given problem. Practical deployment typically involves running active testing only once, without knowledge of the true risk or the ability to rerun the approach multiple times. To help gauge performance in this setting, we derive a bootstrap estimator [7] of the risk-estimation error. Experimentally our approximate 95% confidence intervals contained the true risk-estimation error 88% of the time, demonstrating their efficacy as a diagnostic tool.
In summary, our work allows for LLMs to be evaluated more accurately and with less labelled test data than current best practices by actively selecting data to label at test-time.
2 Background
Our aim is to evaluate a fixed target model, $f$, mapping inputs $x\in\mathcal{X}$ to outputs $y\in\mathcal{Y}$. We formalise evaluation as estimating the frequentist risk of $f$ [3], i.e. the expected loss $R=\mathbb{E}_{p_{\mathrm{eval}}(x,y)}\!\left[\ell(f(x),y)\right]$, where $p_{\mathrm{eval}}$ represents a reference system used as a source of ground truth, and $\ell$ denotes a loss function. As in past work on active testing [16, 17], we consider a pool-based setting [20] where we have access to a pool of $N$ unlabelled test inputs, $\mathcal{D}_{\mathrm{pool}}=\{x_{i}\}_{i=1}^{N}$, but acquiring a label, $y\sim p_{\mathrm{eval}}(y|x)$, for any given $x$ is costly, so we can only afford to acquire $M<N$ labels.
Uniform-random sampling A straightforward way to estimate $R$ is to sample $M$ inputs uniformly at random from the pool and compute $\hat{R}_{\mathrm{unif}}=\frac{1}{M}\sum_{m=1}^{M}\ell(f(x_{i_{m}}),y_{i_{m}})$, where $i_{1:M}\sim\mathrm{Uniform}(\{1,2,\ldots,N\},M)$ are samples drawn without replacement. If the pool was constructed by sampling $x_{i}\sim p_{\mathrm{eval}}(x)$, then $\hat{R}_{\mathrm{unif}}$ is known as the subsample empirical risk and is an unbiased estimator of $R$. However, it will typically have high variance for small values of $M$.
Sampling-based active testing Active testing deals with how to more carefully select the $M$ inputs for labelling to produce a more accurate estimate of $R$. In general this can be done in either a sampling-based or interpolation-based manner. The sampling-based approach uses Monte Carlo estimators of $R$, similar to $R_{\mathrm{unif}}$ but with acquisition indices sampled from a non-uniform distribution over the pool indices. Namely, at acquisition step $m$ we sample pool input $x_{i}$ with probability proportional to $a_{m}(x_{i})\in\mathbb{R}^{+}$, $q_{m}(i)=\frac{a_{m}(x_{i})}{\sum_{x_{j}\in\mathcal{D}_{\mathrm{pool}}}a_{m}(x_{j})}$, where $a_{m}$ is an acquisition function that measures a notion of the expected value of acquiring a label for a given input (the $m$ subscript denotes that it can depend on all the data available at step $m$, including the target model’s training data).
A naive estimator of the risk is $\hat{R}_{\mathrm{naive}}=\frac{1}{M}\sum_{m=1}^{M}\ell(f(x_{i_{m}}),y_{i_{m}})$, where $i_{m}\sim q_{m}(i)$. Farquhar et al, [8] showed that $\hat{R}_{\mathrm{naive}}$ is biased and introduced the levelled unbiased risk estimator (LURE):
$\hat{R}_{\mathrm{LURE}}=\frac{1}{M}\sum_{m=1}^{M}v_{m}\ell(f(x_{i_{m}}),y_{i_{m}}),\quad v_{m}=1+\frac{N-M}{N-m}\left(\frac{1}{(N-m+1)q_{m}(i_{m})}-1\right),$ (1)
where $i_{m}\sim q_{m}(i)$. Given both $\hat{R}_{\mathrm{unif}}$ and $\hat{R}_{\mathrm{LURE}}$ are unbiased, any advantage that $\hat{R}_{\mathrm{LURE}}$ brings comes from variance reduction through a well-designed acquisition function, $a_{m}$.
The optimal acquisition function is $a^{*}_{m}(x)=\mathbb{E}_{p_{\mathrm{eval}}(y|x)}\!\left[\ell(f(x),y)\right]$, i.e. the expected loss under $p_{\mathrm{eval}}(y|x)$. Since this is unknown, Kossen et al, [16] proposed approximating $p_{\mathrm{eval}}(y|x)$ with a surrogate model, $\pi_{m}(y|x)$, which is typically trained on the acquired test labels along with the target model’s training data. This then allows us to define a practical acquisition function:
$a_{m}(x)=\mathbb{E}_{\pi_{m}(y|x)}\!\left[\ell(f(x),y)\right].$ (2)
If $f(x)=p_{f}(y|x)$ and $\ell(\hat{p},y)=-\log\hat{p}(y)$ then the acquisition function becomes $a_{m}(x)=\mathrm{H}\!\left[\pi_{m}(y|x)\,\|\,p_{f}(y|x)\right]$, the cross entropy between the surrogate and target models. Intuitively we can understand this acquisition function as measuring the disagreement between the surrogate and target models, leading us to acquire labels for the inputs where the two models’ predictions differ the most.
Algorithm 1 Sampling-based active testing
0: Target model, $f$; loss function, $\ell$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; budget, $M$
1: Compute $f(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
2: Set $\mathcal{D}_{\mathrm{test}}=\emptyset$
3:for $m\in(1,2,\ldots,M)$ do
4:  Train $\pi_{m}$ (e.g., on $\mathcal{D}_{\mathrm{train}}\cup\mathcal{D}_{\mathrm{test}}$)
5:  Compute $a_{m}(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
6:  Sample $i_{m}\sim q_{m}(i)$
7:  Sample $y_{i_{m}}\sim p_{\mathrm{eval}}(y|x_{i_{m}})$
8:  Compute $\ell(f(x_{i_{m}}),y_{i_{m}})$ and $v_{m}$
9:  Set $\mathcal{D}_{\mathrm{test}}\leftarrow\mathcal{D}_{\mathrm{test}}\cup\{(x_{i_{m}},y_{i_{m}})\}$
10:end for
11: Compute $\hat{R}_{\mathrm{LURE}}$
11: Risk estimate, $\hat{R}_{\mathrm{LURE}}$
Algorithm 2 Interpolation-based active testing
0: Target model, $f$; loss function, $\ell$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; budget, $M$; acquisition function, $a$
1: Compute $f(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
2: Set $\mathcal{D}_{\mathrm{test}}=\emptyset$.
3:for $m\in(1,2,\ldots,M)$ do
4:  Train $\pi_{m}$ (e.g., on $\mathcal{D}_{\mathrm{train}}\cup\mathcal{D}_{\mathrm{test}}$)
5:  Compute $a(x_{j},\pi_{m})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
6:  Select $i_{m}=\arg\max_{j}a(x_{j},\pi_{m})$
7:  Sample $y_{i_{m}}\sim p_{\mathrm{eval}}(y|x_{i_{m}})$
8:  Set $\mathcal{D}_{\mathrm{test}}\leftarrow\mathcal{D}_{\mathrm{test}}\cup\{(x_{i_{m}},y_{i_{m}})\}$
9:end for
10: Compute $\hat{R}_{\mathrm{ASE}}$
10: Risk estimate, $\hat{R}_{\mathrm{ASE}}$
Interpolation-based active testing Kossen et al, [17] introduced an alternative approach to active testing that, like the sampling-based approach, uses a surrogate model to guide data acquisition, but also uses it to estimate the risk. Their active surrogate estimator (ASE) is
$\hat{R}_{\mathrm{ASE}}=\frac{1}{N}\sum\nolimits_{x_{i}\in\mathcal{D}_{\mathrm{pool}}}\mathbb{E}_{\pi_{m}(y|x_{i})}[\ell(f(x_{i}),y)].$ (3)
Here the risk is not directly estimated using labels acquired from $y\sim p_{\mathrm{eval}}(y|x)$ but instead using labels simulated from the surrogate model. The goal of data acquisition is then to improve the surrogate model in a way that leads to a more accurate estimate of the risk. Kossen et al, [17] approached this using an acquisition function called the expected weighted disagreement (XWED).
3 Scaling Up Active Testing
With an understanding of standard approaches, we can now address barriers to scaling up active testing to LLMs. We focus on three key computational bottlenecks: training the surrogate model, computing predictions with the surrogate model, and computing predictions with the target model. Our aim is to reduce the cost of these steps while maintaining the efficacy of active testing.
3.1 Surrogate-model training
We argue that the top priority for reducing the cost of data acquisition is to rethink the training of the surrogate model. Existing approaches involve repeatedly running gradient-based training on the data acquired during testing, typically combined with the target model’s training data. This can be very expensive, especially when working with large datasets and surrogates, as we are focusing on here.
To avoid this expense, we suggest a stripped-back approach: construct the surrogate model using in-context learning [4] on a small amount of randomly acquired data, and fix the surrogate model thereafter. This reduces the cost of surrogate-model training to essentially the minimum possible (other than no training at all). This design decision has important implications for the relative merits of sampling-based ($\hat{R}_{\mathrm{LURE}}$) and interpolation-based ($\hat{R}_{\mathrm{ASE}}$) active testing. We are choosing to use a relatively crude surrogate model to reduce computational cost. This affects $\hat{R}_{\mathrm{LURE}}$ solely through the data-acquisition step (and related corrective weights $v_{m}$), whereas it affects $\hat{R}_{\mathrm{ASE}}$ not only through data acquisition but also much more directly through the expectation over $\pi_{m}(y|x)$ in Equation˜3. Due to this difference in sensitivity between the approaches, we choose $\hat{R}_{\mathrm{LURE}}$ as our default risk estimator for this choice of surrogate model, and demonstrate in Section˜5 that this is well-justified empirically.
3.2 Surrogate-model predictions
Another cost is that of computing predictions with the surrogate model. In sampling-based active testing these predictions arise when computing the acquisition distribution, $q_{m}(i)$, and in interpolation-based active testing they additionally arise when computing the risk estimate, $\hat{R}_{\mathrm{ASE}}$. Our choice to use a fixed surrogate model, $\pi_{0}$, automatically leads to a significant reduction in how many of these predictions are required: we only need to compute $\pi_{0}(y|x_{j})$ once for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$. On top of this we can make each individual forward pass cheaper by using a smaller surrogate model.
3.3 Target-model predictions
The third cost we aim to reduce is that of computing predictions with the target model. The standard versions of both sampling-based and interpolation-based active testing require computing $f(x_{j})$ for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$, which is especially expensive for large target models and pool sets.
This is unavoidable in the interpolation-based approach due to the construction of $\hat{R}_{\mathrm{ASE}}$. But in the sampling-based approach there is scope for an approximation that can reduce the number of target-model predictions from $N$ to $M$, where often $M\ll N$. In particular, when computing the acquisition distribution, $q_{m}(i)$, we can use the surrogate model to approximate the target model’s predictions on the pool set. This leads to a new acquisition function,
$\hat{a}_{m}(x)=\mathbb{E}_{\pi_{m}(y|x)}\!\left[\ell(\psi[\pi_{m}(\cdot|x)],y)\right],$
where $\psi$ denotes an operation that accounts for the fact that the first argument of $\ell$ might not be a probability distribution (e.g., if $\ell(\hat{y},y)=\|\hat{y}-y\|_{2}^{2}$, then we require something like $\psi[\pi_{m}(\cdot|x)]=\mathbb{E}_{\hat{y}\sim\pi_{m}(\cdot|x)}\!\left[\hat{y}\right$]). This acquisition function can be seen as an approximation to $a_{m}(x)$ in Equation˜2. If $\ell(\hat{p},y)=-\log\hat{p}(y)$ then $\hat{a}_{m}(x)=\mathrm{H}\!\left[\pi_{m}(y|x)\right]$, the predictive entropy of the surrogate model.
3.4 Active testing for dataset curation
The use of the surrogate model to approximate the target model during data acquisition also makes it possible to use active testing for dataset curation, i.e. when all the inputs in the pool set are already labelled and the goal is to reduce the number of target-model predictions required for evaluation [23, 28, 29, 35]. This can be viewed as a variation on standard active testing where here we want to use the surrogate model to select $M<N$ input-label pairs with which to evaluate the target model. The acquisition function in this case can depend on the labels as well as the inputs:
$a_{m}(x,y)=\ell(\psi[\pi_{m}(y|x)],y).$ (4)
If we use the logarithmic loss as before, this recovers the negative log likelihood of the surrogate.
4 Estimating Risk-Estimation Error
Aside from its scalability, active testing poses a practical challenge in assessing the quality of the risk estimates it produces. In particular, while the accuracy of simple Monte Carlo error estimates can easily be assessed through calculating the standard error across datapoints, previous work on active testing has not provided any comparable single run error estimator. Given practical deployment of active testing by definition entails only a single run, this leaves practitioners unaware of how well their implementation is working, making it harder to justify active testing’s use.
We address this by proposing a novel estimator of the mean squared error (over possible sequences of acquired indices, $i_{1:K}\sim q(i_{1:K})$, where $K\leq M$) of $\hat{R}_{\mathrm{LURE}}$. For a generic risk estimator, $\hat{R}(i_{1:K})$, we can decompose the mean squared error into bias and variance:
[…]
Letting $L=\{v_{m}\ell(f(x_{i_{m}}),y_{i_{m}})\}_{m=1}^{K}$ denote the sequence of precomputed reweighted losses (as used in Equation˜1), a bootstrap estimator [7] of the risk, $R$, is $\hat{R}_{\mathrm{boot}}=\frac{1}{K}\sum_{m=1}^{K}L_{j_{m}}$, where $j_{m}\sim\mathrm{Uniform}(1,2,\ldots,K)$. Now if we have $B$ bootstrap estimates of the risk, $(\hat{R}_{\mathrm{boot}}^{b})_{b=1}^{B}$, we can form a bootstrap estimator of $\textsc{Var}(\hat{R}_{\mathrm{LURE}})=\mathrm{MSE}(\hat{R}_{\mathrm{LURE}})$:
$\widehat{\textsc{Var}}_{\mathrm{boot}}(\hat{R}_{\mathrm{LURE}})=\frac{1}{B-1}\sum_{b=1}^{B}(\hat{R}_{\mathrm{boot}}^{b}-\bar{R}_{\mathrm{boot}})^{2},$ (6)
where $\bar{R}_{\mathrm{boot}}$ is the mean of the bootstrap estimates, namely $\bar{R}_{\mathrm{boot}}=\frac{1}{B}\sum_{b=1}^{B}\hat{R}_{\mathrm{boot}}^{b}$. We note that $\hat{R}_{\mathrm{boot}}$ is not a standard bootstrap estimator because the reweighted losses, $L_{j}$, are not independent and identically distributed (they are the result of active data acquisition), and the bootstrap resampling process does not capture the dependencies between the losses. As a result, $\hat{R}_{\mathrm{boot}}$ lacks theoretical convergence guarantees. Nevertheless, in Section˜5.9 we find it to be reliably accurate for our experiments, and thus that it provides useful estimates.
5 Experiments
We now seek to empirically assess our proposed active testing approach for LLMs, as well as our proposed risk-error-estimation method (Section˜4). We provide implementation details in Appendix˜A and code for running our experiments at anonymous.4open.science/r/scaling-up-active-testing-0.
5.1 Setup
We focus our evaluation on the following five well-established text classification tasks: Stanford Sentiment Treebank 2 (SST-2; 31), Subjectivity (Subj; 26), Financial Phrase Bank (FPB; 22), Hate Speech (HS; 6), and Massive Multitask Language Understanding (MMLU; 36).
The models we use are the 7B and 70B versions of Llama-2 [34] and Gemma-3 4B [33]. We use these models with either zero- or few-shot prompting. For Llama, we denote the four configurations as 7B-zero, 70B-zero, 7B-few and 70B-few; for Gemma, we use Gemma-3 4B-zero and Gemma-3 4B-few. We always include appropriate task prompts such as “Classify the sentiment of the following sentence as positive or negative”. When we use a few-shot surrogate with $n$ labelled in-context examples to evaluate a zero-shot target, we ensure a fair comparison with uniform sampling by comparing our results using $x$ acquired labels to random sampling results using $x+n$ acquired labels.
Figure 2: Cheap surrogate models support effective data acquisition for active testing. We compare uniform-random testing to active testing (LURE) across four datasets and four target-model configurations (7B-zero, 70B-zero, 7B-few and 70B-few). To guide active data acquisition we use a surrogate (Llama2-7B/-70B in (a), Gemma3-4B in (b)) that we train using a single step of in-context learning and then fix. This stripped-back surrogate-model training, along with the use of small surrogate models relative to the target models, allows us to drastically reduce the computational cost of active testing while achieving strong performance.
We use a logarithmic loss, as discussed in Sections˜2 and 3. Like in past work [16, 17], we measure the performance of a testing method using the median squared error of its risk estimate across random seeds, where the ground-truth risk is computed using a held-out test set.
5.2 Fixed surrogate models trained by in-context learning support effective data acquisition
First we investigate whether our switch from gradient-based training of the surrogate model to a single step of in-context learning still allows accurate risk estimation. We do this by running data acquisition across four datasets, four target models and three testing methods. In particular we compare uniform-random testing against two versions of active testing, LURE-7B and LURE-70B, differing only in the surrogate model used (7B-few vs 70B-few).
To show the benefits of active testing for risk estimation more clearly, we plot the active risk-estimation error as a ratio of the uniform-random risk-estimation error, which we refer to as the relative error. We find that active testing performs very well relative to the standard uniform-random testing in the vast majority of settings (Figure˜2(a)), with risk-estimation error lowered by 32% on average (median over all the experimental variables listed above, as well as over possible label budgets between 1 and 400). This is an exciting result: a practically straightforward technique allows us to improve on standard practice for LLM evaluation with minimal computational overhead. Contrasting with the overall positive result, LURE-70B performs poorly in evaluating the 70B-few model on the SST-2 dataset. We explore this failure case in Section˜5.8 and show it is due to a large number of incorrect labels in the dataset.
Figure 3: Models can both be targets and surrogates in active testing. Although Gemma-3 is smaller than the Llama family, its capabilities are comparable. Llama models—including the 7B variant—can effectively evaluate it, demonstrating that older models can serve as reliable surrogates for newer ones.
5.3 Small surrogate models can be used to evaluate much larger target models
A key additional result to highlight in Figure˜2(a) is that LURE-7B works well not just for evaluating a target model of the same size, but also for evaluating a target model ten times larger. In fact, we show that we can reduce surrogate costs even further by using Gemma-3 4B (Figure˜2(b), bottom row) and Phi-2 (Figure˜C.1 in the Appendix) to evaluate Llama models, reducing computational costs by using even smaller surrogates. Conversely, Llama models effectively evaluate Gemma-3, whose capabilities rival those of larger Llama models on these datasets (Tables C.7 and C.8). As discussed in Section˜3.2, this ability to use a small surrogate model—combined with the fact that fixing the surrogate model allows us to only compute one forward pass on the pool set—is a useful way of reducing costs.
5.4 Using stripped-back surrogate models favours sampling-based active testing
Next we evaluate our claim in Section˜3.1 that sampling-based active testing is preferable over interpolation-based active testing given that the surrogate model is fixed. We do this by comparing our approach with ASE in evaluating the 7B-few and 70B-few models. Our results show ASE produces strong results on the HS dataset with the 70B-few model, but otherwise significantly underperforms (Figure˜4). This is consistent with our explanation that ASE is more sensitive to the quality of the surrogate model, and that computational constraints on the surrogate model make it less practically useful. Note that the ASE risk estimator is constant if the surrogate model is fixed, as it is here.
Figure 4: Our sampling-based active testing (LURE) approach works more reliably than interpolation-based active testing (ASE) when using cheaply constructed surrogate models. In ASE, the surrogate model not only guides data acquisition but also much more directly affects the risk estimate: the labels used to compute the expected loss of the target model are simulated from the surrogate model.
5.5 Good data acquisition is possible without computing predictions with the target model
Now we explore our suggestion in Section˜3.3 that the surrogate model can be used to approximate the target model to completely avoid the need for any target model evaluations during data acquisition. Since the loss is logarithmic, this equates to using the predictive entropy of the surrogate model as the acquisition function. We assess this approach as applied to evaluating the 70B-few model with LURE-7B. We find using the surrogate model to approximate the target model during data acquisition is surprisingly effective (Figure˜5(a)). This finding is very promising: it suggests all of the cost-reduction measures we proposed in Section˜3 are compatible with strong active-testing performance.
Figure 5: (5(a)) Approximating the target model with the surrogate model during data acquisition (here this means acquiring based on the predictive entropy of the surrogate model) reduces the need to compute predictions with the target model while still providing good data acquisition. (5(b)) Active testing with a label-aware acquisition function (here the negative log likelihood of the surrogate model) enables dataset curation tailored to a target model of interest. In the curation problem setting we have access to labels for all inputs in the pool set and the goal is to select a subset of them to reduce the number of target-model predictions required to estimate its risk.
Figure 6: Active testing on the challenging MMLU task remains efficient in most cases, with graceful failure in hardest settings. Notably, Gemma-3 4B is an efficient surrogate, despite belonging to a different model family.
5.6 Active testing for dataset curation is viable
Here we consider the alternative problem setting discussed in Section˜3.4, where the inputs in the pool set are already labelled but it is too expensive to compute target-model predictions on all of them, with a possible solution being to curate a subset using a cheap surrogate model. More concretely we run active testing using the negative log likelihood as our label-aware acquisition function (Equation˜4). We find that our approach improves over uniform-random subsampling on three datasets out of four, and roughly matches on FPB (Figure˜5(b)). These results—along with those for the predictive-entropy function we considered in Section˜5.5—suggest that active testing could be a useful tool for dataset curation, reducing computational costs when using existing evaluation datasets. The practical benefits of this could be significant: Liang et al, [21] required almost 20,000 hours of (Nvidia A100) GPU hours to evaluate 30 models on their HELM benchmark.
Figure 7: The failure of LURE-70B in efficiently evaluating the 70B-few model on the SST-2 dataset (left) is resolved by filtering incorrectly labelled datapoints from the pool set (middle, right). We filter points with a high loss (NLL) and manually confirm that most labels for these points are incorrect (see main text).
5.7 Active evaluation on harder tasks: the MMLU dataset
We next assess the robustness of active testing on the more challenging MMLU dataset [36], where model—and thus surrogate—performance degrades substantially (Tables C.7 and C.8), making data acquisition harder. Despite this, active testing continues to outperform uniform sampling, with graceful failures in the hardest evaluation setting, 70B-few in (Figure˜6). As task difficulty increases, weaker surrogates become insufficient but fail gracefully, as discussed in Section˜C.2. Interestingly, Gemma-3 4B is particularly efficient to evaluate Llama models, highlighting the benefits of diversity when choosing a surrogate from a different model family than the target for cross-entropy sampling. These results show that even small, sufficiently capable models can enable effective active testing.
5.8 Incorrect labels can cause problems
Now we investigate the failure of LURE-70B to effectively evaluate the 70B-few model on the SST-2 dataset (Section˜5.2). We construct two modified versions of the SST-2 pool set: one filtered to remove inputs for which 70B-few’s negative log likelihood (NLL) is greater than 5, and one filtered with a NLL threshold of 3. Then we re-run active testing with these modified pool sets.
We show that the failure case is reduced by the weaker filtering and completely resolved by the stronger filtering (Figure˜7). To understand why, we inspect the filtered-out data and find it is often mislabelled. We find 75% of the 45 examples with NLL greater than 5 are mislabelled, and 67% of the 136 examples with NLL greater than 3. This is a useful demonstration of how active testing can fail when the source of labels is unreliable. Interestingly it can also be seen as a case for using less powerful surrogate models to guide data acquisition: LURE-7B worked well in the exact setting where LURE-70B failed. It is also worth noting that in a dataset-curation setting (Section˜5.6) we have access to all the test labels and therefore have the option to directly filter out high-NLL examples.
Figure 8: Our bootstrap estimator of the risk-estimation error in active testing provides a useful indicator of performance within a given data-acquisition rollout. High coverage probabilities mean the confidence interval is very likely to contain the true risk-estimation error.
5.9 Bootstrap estimation of risk-estimation error provides an accurate performance indicator
Having assessed our proposed changes to active testing, we turn to our proposed method for estimating the mean squared error of the LURE risk estimator (Section˜4). We run active testing $T=100$ times, generating 100 sequences of acquired indices, $(i_{1:M})_{j=1}^{100}$. Then for $K\in\{1,2,\ldots,M\}$ we perform four steps. First, we compute a ground-truth mean-squared error, $\mathrm{MSE}_{K}$, using Equation˜5 with $q(i_{1:K})$ defined as a uniform distribution over the $K$ index sequences acquired so far. Second, we compute a boostrap estimate of the MSE, $\widehat{\mathrm{MSE}}_{K}$, using Equation˜6 with $B=1000$ bootstrap risk estimates. Third, we compute the MSE-estimation error, $(\mathrm{MSE}_{K}-\widehat{\mathrm{MSE}}_{K})^{2}$. Fourth, we construct an approximate 95% confidence interval, $\widehat{\mathrm{MSE}}_{K}\pm 2\hat{\sigma}$, where $\hat{\sigma}$ is the estimated standard deviation.
Our results show MSE-estimation error converging to low values for most runs, after an initial period ($K<100$) of relatively high error (Figure˜8). In addition to this, our confidence intervals provide high levels of coverage: for a strong majority of runs we see coverage probabilities of around 94% for $K\geq 100$, suggesting a practitioner can expect their computed confidence interval to contain the true mean squared error approximately 94% of the time. While there are cases where MSE estimation fails, the overall performance is promising, suggesting our proposed method could be a useful practical performance indicator, helping justify real-world deployment of active testing.
6 Related Work
The sampling-based approach to active testing that our approach builds on was proposed by Kossen et al, [16], using risk estimators from Farquhar et al, [8]. Earlier work on active testing proposed alternative methods. Bennett & Carvalho, [2], Ji et al, [14], Katariya et al, [15] and Kumar & Raj, [19] explored stratification-based techniques, such as splitting the pool set into strata by model confidence and sampling uniformly within each stratum. Sawade et al, [30] and Yilmaz et al, [37] proposed importance-sampling and Poisson-sampling methods related to that of Kossen et al, [16] but without the latter’s innovations of adaptive, surrogate-based data acquisition and risk estimation that accounts for sampling without replacement. Nguyen et al, [24] studied a special case of active testing focused on human vetting of noisy labels.
Recent years have seen various extensions of this earlier active-testing work. Huang et al, [12] introduced a clustering-based approach that relies on computing predictions through the target model on the pool set, which we have argued is a barrier to scaling up. Su et al, [32] proposed a technique tailored to “dense” recognition tasks in computer vision (e.g., image segmentation and object detection). Yu et al, [38] incorporated active testing into the process of training a model. Ashury-Tahan et al, [1] and Hara et al, [10] proposed methods for model selection.
The idea of reducing model-evaluation costs by using carefully selected test datasets has been highlighted a number of times in recent work. Maynez et al, [23] found that the datasets they studied could be substantially reduced in size, through uniform-random subsampling, while maintaining stable rankings of the models they were comparing. Polo et al, [28], Saranathan et al, [29] and Vivek et al, [35] demonstrated more sophisticated methods for achieving the same goal. The goal in these studies was to take an existing test dataset (with labels for all inputs) and reduce it in a way that supports comparisons between different models. This prior work is therefore complementary to our contribution: we primarily study the problem of acquiring new data with which to evaluate a model of interest; and we use a method that tailors data acquisition to that given model.
7 Conclusion
We have argued that a significant barrier to principled active testing of large models is the computational cost of deciding which labels to acquire. To address this we have identified key contributors to the cost—training the surrogate model, and computing predictions with the surrogate and target models—and have proposed straightforward cost-saving measures. We have shown that these measures are compatible with effective active testing, producing low-error risk estimates for large language models. On top of this we have demonstrated an estimator of risk-estimation accuracy that can provide an on-the-fly indication of how well active testing is working on a given practical problem. Overall we believe this represents substantial progress towards more efficient evaluation of LLMs, with significant empirical improvements over current standard practice demonstrated.
Appendix A Experiment Details
A.1 Datasets
Our experiments are based on four datasets: Stanford Sentiment Treebank 2 (SST2) ([31], unknown license), Subjectivity (Subj) ([26], Creative Commons Attribution 4.0 International) Financial Phrase-bank (FPB) ([22], Creative Commons Attribution Non Commercial Share Alike 3.0 Unported License), Hate Speech (HS) ([6], Creative Commons Attribution-Share Alike 3.0 Spain License) and Massive Multitask Language Understanding (MMLU) ([11], MIT license).
Each method, including uniform sampling, samples from the same active set, although its size may vary between experiments. We compare error estimates with the error of a separate test dataset, except for the FPB dataset for which we use the error of the active set due to its reduced size. Table˜A.1 summarizes sizes of the active set and test set for each dataset.
Table A.1: Active set and test set sizes for each dataset. Sets are randomly sampled so that they are disjoints.
SST-2 Subj FPB HS MMLU
Active set 10,000 6,000 2,200 6,000 7,000
Test set 10,000 4,000 4,000 7,000
Stanford Sentiment Treebank 2 (SST-2)
[…]
Subjectivity (Subj)
The Subjectivity dataset is composed of 55k subjective sentences from movie reviews and 55k objective sentences from plot summaries. This dataset is used to determine sentiment polarity. The task is to classify each sentence as either ‘subjective’ or ‘objective’. This dataset is balanced, with $49.5\%$ of objective and $50.5\%$ of subjective sentences. We randomly split the 88k-sentence train set into a 66k-active set and a 44k-test set. Examples for in-context-learning are selected from the separate 22k-sentence test set.
[…]
Financial Phrase-bank (FPB)
[…]
Hate Speech (HS)
[…]
Massive Multitask Language Understanding (MMLU)
[…]
A.2 Models
We employ LLMs from the Llama-2 family ([34], Llama 2 Community License). These are the 7B and 70B versions. We use 8 bit-quantisation for the Llama-2-7B model and half-precision floating-point numbers for the Llama-2-70B model. Kossen et al, [18] showed that these approximations do not significantly affect performance. We additionally employ Gemma-3 4B ([33], Gemma License).
A.3 Prompt formatting
[…]
A.4 Evaluation
For each input, the model outputs deterministic scores for each token in the vocabulary. We select the tokens corresponding to the labels of the dataset and apply the softmax function to obtain a final probability distribution over the possible labels. Any token that does not correspond to one of the labels is ignored.
A.5 Token selection
Tokenizers vary between models, and ideally, one token should correspond to one label, but this is not always the case. As highlighted by Kossen et al, [18], some models might represent task labels with multiple tokens. In such cases, we select only the relevant token that corresponds to the core of the word. For instance, the LLaMa-2 tokenizer encodes the word ‘objective’ as [12091] but ‘subjective’ as [4967, 573], which are tokens for ‘subject’ and ‘ive’. In this case, we select [12091] for ‘objective’ and [4967] for ‘subjective’.
A.6 In-context examples
[…]
A.7 Sampling Details
To obtain deterministic, reproducible token generation, we set the maximum number of tokens to $1$, do not set a top-k nor top-p value and output the logits directly, from which we compute probabilities. As done in Kossen et al, [16], we clip values of the acquisition function to ensure that there is no zero-mass point and thus guarantee that the LURE is unbiased. In practice, all probabilities below $\alpha$ times the probability obtained by a uniform acquisition strategy are brought up to this limit value, where we set $\alpha=0.1$.
A.8 Computational resources
Our experiments are designed to be computationally efficient and therefore do not require substantial resources. Generating model predictions over the pool set is the main computational cost, typically requiring two GPUs for Llama-2 70B and one GPU for all other models used in this paper. Aside from this, the active testing procedure itself can be run efficiently on a small number of CPUs.
Appendix B Impact Statement
This paper presents work whose goal is to generally advance label- and compute-efficiency in model evaluation. As this is not limited to a particular area of application, it is difficult to evaluate the impact of this work. Generally speaking, we believe our work could play a role in reducing the computational costs, and thus the CO2 footprint, associated with model evaluation. As is often the case with research, we would caution against careless application of our method to real-world applications.
Appendix C Additional results
C.1 Active evaluation of Llama models with Phi-2 and Gemma-3 1B surrogates
We further test active evaluation using even smaller surrogates: Phi-2 (2.7B, [13], MIT License) in Figure˜C.1 and Gemma-3 1B ([33], Gemma License) in Figure˜C.2. These models are limited in their capabilities. Although they generally perform well when evaluating 70b-zero or 7B-zero, they struggle with more performant few-shot configurations, leading to graceful failures comparable to random sampling. These results illustrate the limits of active evaluation with weak surrogates: effective active sampling depends on the surrogate providing informative signal, and when it does not, active testing offers little to no advantage over uniform sampling.
Figure C.1: Even very small surrogate models, here Phi-2, can achieve label-efficient evaluation in simple cases, although they struggle with evaluating our best performing target model, 70B-few.
Figure C.2: Even very small surrogate models, here Gemma-3 1B, can achieve label-efficient evaluation in simple cases, although they struggle with evaluating our best performing target model, 70B-few.
C.2 Weak surrogates perform worse on the more challenging MMLU dataset
[…]
Figure C.3: On the harder MMLU task, weak surrogates become insufficient. Yet, active evaluation with an uninformative surrogate fails gracefully, with performance close to uniform sampling and similar risk-estimation error decrease. This demonstrates the robustness of active testing in inefficient settings.
C.3 Examination of Pearson’s correlation coefficient
While we previously examined the failure of active evaluation on SST-2 with LURE-70B for the 70B-few configuration (Figure˜2), we also observe underperformance on Subj when using LURE-7B to evaluate 70B-few. Unlike the SST-2 case, filtering out high-loss points does not improve performance. To better understand this, we analyze in Table˜C.1 the Pearson correlation coefficient between the cross-entropy-based acquisition scores and the optimal acquisition function (negative-log likelihood of the target’s predictions). This correlation quantifies how well the surrogate approximates the optimal acquisition strategy. Notably, this is the only setting where the correlation is close to zero, and negative–indicating that the surrogate provides no meaningful guidance and the corresponding acquisition proposal may be selecting examples essentially at random. This explains why active evaluation performs similarly to uniform sampling in this case.
Table C.1: Pearson’s correlation coefficient between the cross-entropy of the surrogate’s and target’s predictions and the negative-log likelihood of the target’s predictions. Low or negative correlation indicates poor alignment between the surrogate and optimal acquisition function, explaining degraded active evaluation performance in specific cases. Active testing failure cases are shown in bold.
Model 70B-few 7B-few
Method LURE-Llama2-70B LURE-Llama2-7B LURE-Llama2-70B LURE-Llama2-7B
FPB $0.628$ $0.392$ $0.707$ $0.588$
SST-2 $\bf{0.285}$ $0.224$ $0.527$ $0.454$
Subj $0.364$ $\bf{-0.025}$ $0.802$ $0.652$
HS $0.461$ $0.146$ $0.555$ $0.384$
C.4 Risk-estimation values
In addition to figures shown in the main paper, we provide risk-estimation errors values for each experiment at acquisition steps $50,100,200,300$, and $400$ in Tables˜C.2, C.3, C.4, C.5 and C.6.
C.5 Evaluation results
Tables˜C.7 and C.8 provide the loss and accuracy numbers for Llama-2 7B, Llama-2 70B, and Gemma-3 4B, in both few- and zero-shot configurations, on our selection of target tasks. Results for Llama-2 here match those published in Kossen et al, [18].
Table C.2: Risk-estimation error for experiments on FPB using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $24.3373$ $11.5663$ $5.1944$ $3.4854$ $2.4122$
LURE-Llama2-7B $8.0511$ $3.6772$ $1.7589$ $1.0463$
LURE-Llama2-70B $\mathbf{4.2225}$ $\mathbf{2.0045}$ $\mathbf{0.98390}$ $\mathbf{0.64870}$
LURE-Gemma3-4B $7.1132$ $3.5176$ $1.6571$ $1.0540$
7B-zero Uniform random $9.4687$ $4.8302$ $2.1823$ $1.4524$ $1.0075$
LURE-Llama2-7B $2.5167$ $1.3094$ $0.59430$ $0.37380$
LURE-Llama2-70B $1.9294$ $0.96850$ $0.42460$ $0.27060$
LURE-Gemma3-4B $\mathbf{1.1421}$ $\mathbf{0.5849}$ $\mathbf{0.3681}$ $\mathbf{0.2341}$
70B-few Uniform random $10.5749$ $5.16310$ $2.48680$ $1.51150$ $0.977800$
LURE-Llama2-7B $6.3984$ $3.2058$ $1.3719$ $0.89110$ $0.63880$
LURE-Llama2-70B $5.0757$ $\mathbf{2.5057}$ $\mathbf{1.1921}$ $\mathbf{0.76740}$ $\mathbf{0.54300}$
LURE-Gemma3-4B $\mathbf{5.0390}$ $2.6267$ $1.3102$ $0.8344$ $0.5825$
7B-few Uniform random $14.3272$ $7.2276$ $3.4608$ $2.3147$ $1.5246$
LURE-Llama2-7B $8.9252$ $4.5853$ $2.3314$ $1.5486$ $1.1345$
LURE-Llama2-70B $\mathbf{4.8174}$ $\mathbf{2.1675}$ $\mathbf{1.1186}$ $\mathbf{0.65290}$ $\mathbf{0.45770}$
LURE-Gemma3-4B $6.5254$ $3.7684$ $1.9489$ $1.2371$ $0.8556$
Table C.3: Risk-estimation error for experiments on SST-2 using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $2.1154$ $1.0992$ $0.56060$ $0.34910$ $0.26150$
LURE-Llama2-7B $0.5386$ $0.3089$ $0.1551$ $0.1023$
LURE-Llama2-70B $\mathbf{0.4191}$ $\mathbf{0.2308}$ $\mathbf{0.1255}$ $\mathbf{0.08190}$
LURE-Gemma3-4B $0.5283$ $0.2790$ $0.1542$ $0.1042$
7B-zero Uniform random $12.4455$ $5.49$ $2.9545$ $1.9694$ $1.362$
LURE-Llama2-7B $2.7160$ $1.4229$ $0.66190$ $0.48410$
LURE-Llama2-70B $\mathbf{2.3740}$ $\mathbf{1.2437}$ $\mathbf{0.60510}$ $\mathbf{0.42530}$
LURE-Gemma3-4B $3.5371$ $1.6843$ $0.8699$ $0.6062$
70B-few Uniform random $36.6017$ $16.8825$ $9.4717$ $6.3474$ $4.3089$
LURE-Llama2-7B $\mathbf{24.8052}$ $\mathbf{14.7678}$ $\mathbf{7.9162}$ $\mathbf{5.2615}$ $\mathbf{4.0596}$
LURE-Llama2-70B $40.8407$ $30.077$ $21.3892$ $17.1808$ $14.7939$
LURE-Gemma3-4B $26.3377$ $16.8812$ $9.4499$ $7.1826$ $5.7175$
7B-few Uniform random $19.5177$ $9.5602$ $4.9347$ $3.1286$ $2.2497$
LURE-Llama2-7B $14.2747$ $8.3696$ $4.6265$ $3.2263$ $2.3764$
LURE-Llama2-70B $\mathbf{9.7216}$ $\mathbf{6.4914}$ $\mathbf{3.6886}$ $2.8971$ $2.1722$
LURE-Gemma3-4B $13.3504$ $7.0426$ $3.9896$ $\mathbf{2.8028}$ $\mathbf{1.9958}$
Table C.4: Risk-estimation error for experiments on Subjectivity using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $1.2104$ $0.64740$ $0.30890$ $0.21120$ $0.15430$
LURE-Llama2-7B $0.4688$ $0.2311$ $0.1113$ $0.07570$
LURE-Llama2-70B $\mathbf{0.139}$ $\mathbf{0.0758}$ $\mathbf{0.0373}$ $\mathbf{0.0251}$
LURE-Gemma3-4B $0.3809$ $0.1897$ $0.0887$ $0.0638$
7B-zero Uniform random $9.7212$ $5.1028$ $2.4378$ $1.5612$ $1.1429$
LURE-Llama2-7B $3.4150$ $1.6835$ $0.87130$ $0.58730$
LURE-Llama2-70B $\mathbf{1.6958}$ $\mathbf{0.89970}$ $\mathbf{0.44150}$ $\mathbf{0.29280}$
LURE-Gemma3-4B $2.8896$ $1.4536$ $0.7144$ $0.4787$
70B-few Uniform random $15.5609$ $6.6315$ $3.7465$ $2.5614$ $2.1347$
LURE-Llama2-7B $16.4075$ $8.2611$ $4.4942$ $3.1802$ $2.1987$
LURE-Llama2-70B $\mathbf{8.2473}$ $\mathbf{5.0366}$ $\mathbf{3.0052}$ $\mathbf{2.1589}$ $\mathbf{1.6293}$
LURE-Gemma3-4B $12.615$ $6.7789$ $3.3058$ $2.2544$ $1.7645$
7B-few Uniform random $7.4461$ $3.9287$ $1.8132$ $1.1308$ $0.87000$
LURE-Llama2-7B $5.2522$ $2.6268$ $1.2944$ $0.84500$ $0.63670$
LURE-Llama2-70B $\mathbf{1.8539}$ $\mathbf{1.0921}$ $\mathbf{0.62250}$ $\mathbf{0.42070}$ $\mathbf{0.30200}$
LURE-Gemma3-4B $4.1246$ $2.1545$ $1.0256$ $0.6868$ $0.4941$
Table C.5: Risk-estimation error for experiments on Hate Speech using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $2.3826$ $1.1493$ $6.3630$ $0.39690$ $0.30590$
LURE-Llama2-7B $0.8497$ $0.4371$ $0.2247$ $0.1524$
LURE-Llama2-70B $\mathbf{0.5703}$ $\mathbf{0.2996}$ $\mathbf{0.1418}$ $\mathbf{0.09900}$
LURE-Gemma3-4B $0.7508$ $0.3700$ $0.1822$ $0.1330$
7B-zero Uniform random $4.3455$ $\mathbf{2.3304}$ $\mathbf{1.3545}$ $0.96420$ $0.80470$
LURE-Llama2-7B $3.8653$ $1.9395$ $1.1326$ $0.86310$
LURE-Llama2-70B $3.0157$ 1.7693{1.7693} $\mathbf{0.94190}$ $\mathbf{0.73000}$
LURE-Gemma3-4B $3.5508$ $1.6981$ $0.9918$ $0.7766$
70B-few Uniform random $24.5497$ $11.8593$ $5.8947$ $4.263$ $2.897$
LURE-Llama2-7B $26.4253$ $13.5746$ $6.9755$ $4.249$ $3.0014$
LURE-Llama2-70B $19.8693$ $11.2704$ $6.1626$ $\mathbf{3.983}$ $2.9785$
LURE-Gemma3-4B $\mathbf{19.4658}$ $\mathbf{10.4853}$ $\mathbf{5.473}$ $\mathbf{3.6247}$ $\mathbf{2.767}$
7B-few Uniform random $23.7706$ $11.7404$ $6.6989$ $4.6215$ $3.4943$
LURE-Llama2-7B $19.6371$ $9.638$ $5.8037$ $4.0226$ $3.1452$
LURE-Llama2-70B $\mathbf{13.4897}$ $\mathbf{7.5196}$ $\mathbf{4.1388}$ $\mathbf{2.7654}$ $\mathbf{2.2691}$
LURE-Gemma3-4B $16.6135$ $8.7585$ $4.8131$ $3.4505$ $2.8044$
Table C.6: Risk-estimation error for experiments on MMLU using the cross-entropy acquisition functions for LURE experiments. All risk-estimation error in the table are scaled by $10^{-4}$. For each experiment, the minimum value is shown in bold.
Acquisition step $50$ $100$ $200$ $300$ $400$
Target Method
70B-zero Uniform random $45.2305$ $22.0405$ $11.5532$ $7.2541$ $5.288$
LURE-Llama2-7B $57.137$ $28.4008$ $13.3008$ $8.4398$ $6.7655$
LURE-Llama2-70B $\mathbf{30.4079}$ 17.3273{17.3273} 9.5205{9.5205} 6.6253{6.6253} $4.9939$
LURE-Gemma3-4B $30.894$ $\mathbf{16.3988}$ $\mathbf{9.0229}$ $\mathbf{5.6838}$ $\mathbf{4.1935}$
7B-zero Uniform random $8.1557$ $4.1386$ $2.1924$ $1.4752$ $1.1799$
LURE-Llama2-7B $6.6835$ $3.2991$ $1.6763$ $1.2791$ $0.9767$
LURE-Llama2-70B $\mathbf{4.3678}$ $\mathbf{2.6164}$ $\mathbf{1.3063}$ $\mathbf{0.9088}$ $\mathbf{0.7330}$
LURE-Gemma3-4B $4.9507$ $2.6543$ $1.4700$ $0.9569$ $0.7622$
70B-few Uniform random 80.9096{80.9096} 41.8724{41.8724} 20.7615{20.7615} 12.5686{12.5686} 8.9944{8.9944}
LURE-Llama2-7B $94.483$ $51.3035$ $27.2732$ $17.5476$ $11.713$
LURE-Llama2-70B $84.9646$ $45.9222$ $23.3469$ $15.7314$ $12.5486$
LURE-Gemma3-4B $\mathbf{71.571}$ $\mathbf{34.8396}$ $\mathbf{19.3527}$ $\mathbf{12.9035}$ $\mathbf{9.2558}$
7B-few Uniform random $20.5838$ $10.142$ $4.8003$ $3.0952$ $2.4424$
LURE-Llama2-7B $16.2533$ $8.686$ $4.5997$ $3.047$ $2.2239$
LURE-Llama2-70B $\mathbf{7.242}$ $\mathbf{4.3427}$ $\mathbf{2.1991}$ $\mathbf{1.4639}$ $\mathbf{1.1166}$
LURE-Gemma3-4B $9.6443$ $4.9703$ $2.3827$ $1.694$ $1.2064$
Table C.7: Loss of the Llama-2 (L) 7B and 70B models and of the Gemma-3 (G) 4B model evaluated on the active set of various datasets.
L 7B-zero L 7B-few L 70B-zero L 70B-few G 4B-zero G 4B-few
FPB $0.7847$ $0.3084$ $1.3123$ $0.2152$ $1.4695$ $0.2701$
SST-2 $0.5639$ $0.2050$ $0.5808$ $0.1828$ $0.6234$ $0.2107$
Subj $0.7326$ $0.3278$ $0.5873$ $0.1235$ $0.7101$ $0.2741$
HS $0.9129$ $0.2979$ $0.5386$ $0.2166$ $0.7245$ $0.2678$
MMLU $1.3468$ $1.2568$ $0.9217$ $0.8384$ $1.0920$ $0.9683$
Table C.8: Accuracy of the Llama-2 (L) 7B and 70B models and of the Gemma-3 (G) 4B model evaluated on the active set of various datasets.
L 7B-zero L 7B-few L 70B-zero L 70B-few G 4B-zero G 4B-few
FPB $66.75$ $90.60$ $26.25$ $94.15$ $29.70$ $91.00$
SST-2 $63.96$ $92.19$ $76.80$ $93.60$ $57.79$ $91.68$
Subj $49.38$ $89.72$ $54.85$ $96.12$ $50.40$ $93.73$
HS $11.38$ $89.28$ $84.18$ $90.23$ $42.87$ $89.42$
MMLU $35.41$ $42.16$ $60.80$ $65.34$ $52.61$ $57.83$