### v2

# Scaling Up Active Testing to Large Language Models

Abstract
Active testing enables label-efficient evaluation of predictive models through careful data acquisition, but it can pose a significant computational cost. We identify cost-saving measures that enable active testing to be scaled up to large language models (LLMs). In particular we show that the surrogate model used to guide data acquisition can be constructed cheaply using in-context learning, does not require updating within an active-testing loop, and can be smaller than the target model. We even find we can make good data-acquisition decisions without making predictions with the target model. As a result we are able to achieve much more accurate evaluations of LLM performance relative to using randomly acquired data. We additionally introduce a bootstrap estimator of evaluation error, which we show to be a useful indicator of how well active testing is working within a single run.
1 Introduction
Evaluating frontier models is becoming more expensive as they become more sophisticated (burden2024evaluating; openai2023gpt4). At the same time, with new models arriving in quick succession, and ever-present scope for data to leak from evaluations to training (ganguli2023challenges), the evaluation problem is a dynamic one: it requires ongoing, adaptive gathering of new evaluation data.
Active testing (kossen2021active; kossen2022active) is an attractive solution to this problem. Motivated by the observation that some labels are more informative than others about the behaviour of a target model, it involves carefully deciding which test inputs to acquire labels for. The foundation for data acquisition is a surrogate model of the test-time label distribution, which is typically updated on the labels acquired during testing. Making a data-acquisition decision requires making predictions with the surrogate model and often also the target model. The computational cost of this process is often justifiable in light of high labelling costs, but it limits the scope of active testing’s applicability.
In this work we show how active testing can be scaled up to large language models (LLMs), allowing us to more effectively evaluate them with limited labelling budgets. We begin by highlighting three computational bottlenecks in the predominant approach to active testing: updating the surrogate model as labels are acquired, making predictions with the surrogate model, and making predictions with the target model. The first of these is particularly important: it has traditionally involved repeatedly running gradient-based training on the acquired test data within an active-testing loop.
Figure 1: Our proposed active-testing approach enables low-error estimates of the risk (expected predictive loss) of a large language model (here the 70B version of Llama 2) on four text-classification problems (SST-2, FPB, HS and Subj) while only using a small label budget. We compare uniform-random testing against active testing (LURE), using either a 7B or 70B surrogate model with in-context learning to guide data acquisition.
We then identify simple and surprisingly effective ways to address each of these key bottlenecks. First, we strip back the training of the surrogate model to a single step of in-context learning (brown2020language) on a small amount of initial test data, removing the need for repeated gradient-based training. Second, we show it is possible to use a surrogate model that is smaller than the target model, reducing the cost of surrogate-model predictions. Third, we show we can even forgo making any predictions with the target model, relying solely on the surrogate model for data acquisition. Together these changes make active testing dramatically cheaper, enabling it to be scaled up to LLMs.
Empirically we find our approach can substantially improve over the common practice of acquiring data uniformly at random (Figure 1). More specifically it can estimate the risk (expected predictive loss over test examples) of LLMs with an estimation error typically 25% to 50%—and sometimes up to 80%—lower than that obtained through uniform-random data acquisition. Our work therefore represents significant progress in accurately and dynamically evaluating LLMs, as well as in more generally increasing the scope for active testing to be used in practical settings.
To further improve the real-world applicability of active testing, we address the challenge of judging how well it is working on a given problem. In research we can assess active testing by running it multiple times and comparing its risk estimates to a known true risk. But practical deployment involves running active testing only once, without knowledge of the true risk, making it hard to gauge active-testing performance. To help address this we derive a bootstrap estimator (efron1979bootstrap) of the risk-estimation error. In experiments we find our approximate 95% confidence intervals contain the true risk-estimation error 88% of the time, suggesting our estimator can be a useful diagnostic tool.
2 Background
Our aim is to evaluate a fixed target model, $f$, mapping inputs $x\in\mathcal{X}$ to outputs $y\in\mathcal{Y}$. We formalise evaluation as estimating a form of frequentist risk (berger1985statistical), namely an expected predictive loss of the form $R=\mathbb{E}_{p_{\mathrm{eval}}(x,y)}\!\left[\ell(f(x),y)\right]$, where $p_{\mathrm{eval}}$ represents a reference system used as a source of ground truth, and $\ell$ denotes a loss function that represents the consequences of predictive errors. As in past work (kossen2021active; kossen2022active), we consider a pool-based setting (lewis1994sequential) where we have access to a pool of $N$ unlabelled test inputs, $\mathcal{D}_{\mathrm{pool}}=\{x_{i}\}_{i=1}^{N}$, but acquiring a label, $y\sim p_{\mathrm{eval}}(y|x)$, for any given $x$ is costly, so we can only afford to acquire $M<N$ labels.
2.1 Uniform-random sampling
A simple estimator of $R$ uses uniform-random samples from the pool:
$\hat{R}_{\mathrm{unif}}=\frac{1}{M}\sum_{m=1}^{M}\ell(f(x_{i_{m}}),y_{i_{m}}),$
where $i_{1:M}\sim\mathrm{Uniform}(\{1,2,\ldots,N\},M)$ are indices sampled without replacement. If the pool was constructed by sampling $x_{i}\sim p_{\mathrm{eval}}(x)$ then $\hat{R}_{\mathrm{unif}}$ is known as the subsample empirical risk and is an unbiased estimator of $R$. However, it will typically have high variance for small values of $M$.
2.2 Sampling-based active testing
Active testing deals with how to more carefully select the $M$ inputs for labelling to produce a more accurate estimate of $R$. This can be achieved through either a sampling-based approach or an interpolation-based approach. In sampling-based active testing we use Monte Carlo estimators of $R$, similar to $R_{\mathrm{unif}}$ but with pool indices sampled from a non-uniform distribution. At acquisition step $m$ we sample pool input $x_{i}$ with probability proportional to $a_{m}(x_{i})\in\mathbb{R}^{+}$ where $a_{m}$ is an acquisition function that measures a notion of how useful we expect the label for a given input to be (the $m$ subscript denotes that it can depend on all the data available at step $m$, including the target model’s training data). That is, we sample $i_{m}\sim q_{m}(i)$ where $q_{m}(i)=a_{m}(x_{i})/\sum_{x_{j}\in\mathcal{D}_{\mathrm{pool}}}a_{m}(x_{j})$.
farquhar2021statistical showed that naive Monte Carlo with $i_{m}\sim q_{m}(i)$ is biased and, to address this, introduced the levelled unbiased risk estimator (LURE):
$\hat{R}_{\mathrm{LURE}}=\frac{1}{M}\sum_{m=1}^{M}v_{m}\ell(f(x_{i_{m}}),y_{i_{m}}),\quad v_{m}=1+\frac{N-M}{N-m}\left(\frac{1}{(N-m+1)q_{m}(i_{m})}-1\right).$ (1)
Given both $\hat{R}_{\mathrm{unif}}$ and $\hat{R}_{\mathrm{LURE}}$ are unbiased, any advantage that $\hat{R}_{\mathrm{LURE}}$ brings comes from variance reduction through a well-designed acquisition function, $a_{m}$. The optimal acquisition function is $a^{*}_{m}(x)=\mathbb{E}_{p_{\mathrm{eval}}(y|x)}\!\left[\ell(f(x),y)\right]$, the expected loss under $p_{\mathrm{eval}}(y|x)$. Since this is unknown, kossen2021active proposed approximating $p_{\mathrm{eval}}(y|x)$ with a surrogate model, $\pi_{m}(y|x)$, which is typically trained on the acquired test labels along with the target model’s training data. This surrogate model then allows us to define a practical acquisition function:
$a_{m}(x)=\mathbb{E}_{\pi_{m}(y|x)}\!\left[\ell(f(x),y)\right].$ (2)
If $f(x)=p_{f}(y|x)$ and $\ell(\hat{p},y)=-\log\hat{p}(y)$ then the acquisition function becomes $a_{m}(x)=\mathrm{H}\!\left[\pi_{m}(y|x)\,\|\,p_{f}(y|x)\right]$, the cross entropy between the surrogate and target models. Intuitively we can understand this acquisition function as measuring the disagreement between the surrogate and target models, leading us to acquire labels for the inputs where the two models’ predictions differ the most.
Algorithm 1 Sampling-based active testing
0: Target model, $f$; loss function, $\ell$; acquisition function, $a$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; label budget, $M$
1: Compute $f(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
2: Set $\mathcal{D}_{\mathrm{test}}=\emptyset$
3:for $m\in(1,2,\ldots,M)$ do
4:  Train $\pi_{m}$ (e.g., on $\mathcal{D}_{\mathrm{train}}\cup\mathcal{D}_{\mathrm{test}}$)
5:  Compute $a_{m}(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
6:  Sample $i_{m}\sim q_{m}(i)$
7:  Sample $y_{i_{m}}\sim p_{\mathrm{eval}}(y|x_{i_{m}})$
8:  Set $\mathcal{D}_{\mathrm{test}}\leftarrow\mathcal{D}_{\mathrm{test}}\cup\{(x_{i_{m}},y_{i_{m}})\}$
9:end for
10: Compute $\hat{R}_{\mathrm{LURE}}$
10: Risk estimate, $\hat{R}_{\mathrm{LURE}}$
Algorithm 2 Interpolation-based active testing
0: Target model, $f$; loss function, $\ell$; acquisition function, $a$; training set, $\mathcal{D}_{\mathrm{train}}$; pool set, $\mathcal{D}_{\mathrm{pool}}$; label budget, $M$
1: Compute $f(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
2: Set $\mathcal{D}_{\mathrm{test}}=\emptyset$
3:for $m\in(1,2,\ldots,M)$ do
4:  Train $\pi_{m}$ (e.g., on $\mathcal{D}_{\mathrm{train}}\cup\mathcal{D}_{\mathrm{test}}$)
5:  Compute $a_{m}(x_{j})$ for all $x_{j}\in\mathcal{D}_{\mathrm{pool}}$
6:  Select $i_{m}=\arg\max_{j}a_{m}(x_{j})$
7:  Sample $y_{i_{m}}\sim p_{\mathrm{eval}}(y|x_{i_{m}})$
8:  Set $\mathcal{D}_{\mathrm{test}}\leftarrow\mathcal{D}_{\mathrm{test}}\cup\{(x_{i_{m}},y_{i_{m}})\}$
9:end for
10: Compute $\hat{R}_{\mathrm{ASE}}$
10: Risk estimate, $\hat{R}_{\mathrm{ASE}}$
2.3 Interpolation-based active testing
kossen2022active introduced an alternative approach to active testing that, like the sampling-based approach, uses a surrogate model to guide data acquisition, but also uses it to estimate the risk. Their active surrogate estimator (ASE) is
$\hat{R}_{\mathrm{ASE}}=\frac{1}{N}\sum\nolimits_{x_{i}\in\mathcal{D}_{\mathrm{pool}}}\mathbb{E}_{\pi_{m}(y|x_{i})}[\ell(f(x_{i}),y)].$ (3)
Here the risk is not directly estimated using labels acquired from $y\sim p_{\mathrm{eval}}(y|x)$ but instead using labels simulated from the surrogate model. The goal of data acquisition is then to improve the surrogate model in a way that leads to a more accurate estimate of the risk. kossen2022active approached this using an acquisition function called the expected weighted disagreement (XWED).
3 Scaling up active testing
With an understanding of standard approaches to active testing, we can now identify and address barriers to scaling up to LLMs. We focus on three key computational bottlenecks: training the surrogate model, making predictions with the surrogate model, and making predictions with the target model. Our aim is to reduce the cost of these steps while maintaining the efficacy of active testing.
3.1 Surrogate-model training
We argue that the top priority for reducing the cost of data acquisition is to rethink the training of the surrogate model. Existing approaches involve repeatedly running gradient-based training on the data acquired during testing, typically combined with the target model’s training data. This can be very expensive, especially when working with large datasets and surrogates, as we are focusing on here.
To avoid this expense we suggest a stripped-back approach: construct the surrogate model using in-context learning (brown2020language) on a small amount of randomly acquired test data, then keep it fixed. This reduces the cost of surrogate-model training to essentially the minimum possible.
This design decision has important implications for the relative merits of sampling-based and interpolation-based active testing. We are using a relatively crude surrogate model to reduce computational cost. This affects the sampling-based estimator solely through the data-acquisition step (and related corrective weights, $v_{m}$), whereas it affects the interpolation-based estimator not only through data acquisition but also much more directly through the expectation over $\pi_{m}(y|x)$ in Equation 3. Due to this difference in sensitivity between the approaches, we default to the sampling-based estimator for this choice of surrogate model, and demonstrate in Section 5 that this is well-justified empirically.
3.2 Surrogate-model predictions
Another cost is that of making predictions with the surrogate model. In sampling-based active testing these predictions arise when computing the acquisition distribution, $q_{m}(i)$, and in interpolation-based active testing they additionally arise when computing the risk estimate.
Our choice to use a fixed surrogate model, $\pi_{0}$, automatically leads to a significant reduction in how many of these predictions are required: we only need to compute $\pi_{0}(y|x_{j})$ once for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$. On top of this we can make each individual forward pass cheaper by using a smaller surrogate model.
3.3 Target-model predictions
The third cost we aim to reduce is that of making predictions with the target model. The standard versions of both sampling-based and interpolation-based active testing require computing $f(x_{j})$ for each $x_{j}\in\mathcal{D}_{\mathrm{pool}}$, which is especially expensive for large target models and pool sets.
This is unavoidable in the interpolation-based approach due to the construction of $\hat{R}_{\mathrm{ASE}}$. But in the sampling-based approach there is scope for an approximation that can reduce the number of target-model predictions from $N$ to $M$, where often $M\ll N$. In particular, when computing the acquisition distribution, $q_{m}(i)$, we can use the surrogate model to approximate the target model’s predictions on the pool set. This leads to a new acquisition function,
$\hat{a}_{m}(x)=\mathbb{E}_{\pi_{m}(y|x)}\!\left[\ell(\psi[\pi_{m}(\cdot|x)],y)\right],$
where $\psi$ denotes an operation that accounts for the fact that the first argument of $\ell$ might not be a probability distribution (e.g., if $\ell(\hat{y},y)=\|\hat{y}-y\|_{2}^{2}$, then we require something like $\psi[\pi_{m}(\cdot|x)]=\mathbb{E}_{\hat{y}\sim\pi_{m}(\cdot|x)}\!\left[\hat{y}\right$]). This acquisition function can be seen as an approximation to $a_{m}(x)$ in Equation 2. If $\ell(\hat{p},y)=-\log\hat{p}(y)$ then $\hat{a}_{m}(x)=\mathrm{H}\!\left[\pi_{m}(y|x)\right]$, the predictive entropy of the surrogate model.
3.4 Active testing for dataset curation
The use of the surrogate model to approximate the target model during data acquisition also makes it possible to use active testing for dataset curation, where all the inputs in the pool set are already labelled and the goal is to reduce the number of target-model predictions required for evaluation (maynez2023benchmarking; polo2024tinybenchmarks; saranathan2024dele; vivek2024anchor). This can be viewed as a variation on standard active testing where here we want to use the surrogate model to select $M<N$ input-label pairs with which to evaluate the target model. The acquisition function in this case can depend on the labels as well as the inputs:
$a_{m}(x,y)=\ell(\psi[\pi_{m}(y|x)],y).$ (4)
If we use the logarithmic loss as before, this recovers the negative log likelihood of the surrogate.
4 Estimating risk-estimation error
Aside from its scalability, active testing poses a practical challenge in assessing the quality of the risk estimates it produces. In research we can construct repeatable simulations of active testing and record ground-truth risk-estimation errors across repeat runs. But practical deployment of active testing involves only a single run, without knowledge of the true risk. This leaves practitioners unaware of how well their implementation is working, making it harder to justify using active testing.
We propose addressing this using a novel estimator of the mean squared error (over possible sequences of acquired indices, $i_{1:K}\sim q(i_{1:K})$, where $K\leq M$) of $\hat{R}_{\mathrm{LURE}}$. For a generic risk estimator, $\hat{R}(i_{1:K})$, we can decompose the mean squared error into squared bias and variance:
$\mathrm{MSE}(\hat{R})=\mathbb{E}_{q(i_{1:K})}[(\hat{R}(i_{1:K})-R)^{2}]=\underbrace{(\mathbb{E}_{q(i_{1:K})}[\hat{R}(i_{1:K})]-R)^{2}}_{\mathrm{Bias}(\hat{R})^{2}}+\underbrace{\mathbb{V}_{q(i_{1:K})}[\hat{R}(i_{1:K})]}_{\mathrm{Var}(\hat{R})}$. (5)
Since $\mathrm{Bias}(\hat{R}_{\mathrm{LURE}})=0$, our task is to estimate $\mathrm{Var}(\hat{R}_{\mathrm{LURE}})=\mathrm{MSE}(\hat{R}_{\mathrm{LURE}})$ from a single sequence of acquired indices, $i_{1:K}$. Letting $L=(v_{m}\ell(f(x_{i_{m}}),y_{i_{m}}))_{m=1}^{K}$ denote the first $K$ reweighted losses used in Equation 1, a bootstrap estimator (efron1979bootstrap) of the risk, $R$, is
$\hat{R}_{\mathrm{boot}}=\frac{1}{K}\sum_{m=1}^{K}L_{j_{m}},\quad j_{m}\sim\mathrm{Uniform}(1,2,\ldots,K).$
If we have $B$ bootstrap estimates, $(\hat{R}_{\mathrm{boot}}^{b})_{b=1}^{B}$, we can form a bootstrap estimator of the variance:
$\widehat{\mathrm{Var}}_{\mathrm{boot}}(\hat{R}_{\mathrm{LURE}})=\frac{1}{B-1}\sum_{b=1}^{B}(\hat{R}_{\mathrm{boot}}^{b}-\bar{R}_{\mathrm{boot}})^{2},\quad\bar{R}_{\mathrm{boot}}=\frac{1}{B}\sum_{b=1}^{B}\hat{R}_{\mathrm{boot}}^{b}.$ (6)
It is important to note that $\hat{R}_{\mathrm{boot}}$ is not a standard bootstrap estimator because the reweighted losses, $L_{j}$, are not independent and identically distributed (they are the result of active data acquisition), and the bootstrap resampling process does not capture the dependencies between the losses. As a result, the estimator lacks theoretical convergence guarantees. Nevertheless we find it to be reliably accurate in our experiments (Section 5.9), suggesting it could be useful in practice.
5 Experiments
We now seek to empirically assess our proposed approach for active testing of LLMs (Section 3), as well as our proposed risk-error-estimation method (Section 4). We provide implementation details in Appendix B and code at github.com/gabrielleberrada/scaling-up-active-testing.
5.1 Setup
We use five text-classification datasets: Stanford Sentiment Treebank 2 (SST-2; socher2013recursive), Subjectivity (Subj; pang2004sentimental), Financial Phrase Bank (FPB; malo2013good), Hate Speech (HS; degibert2018hate) and Massive Multitask Language Understanding (MMLU; hendrycks2021measuring). We mainly focus on the first four, which were previously used by kossen2024context, then use MMLU to explore a problem that is more challenging for the models used in our experiments.
The models we use span three model families and have parameter counts ranging from 1 to 70 billion. In our main experiments we use the 7B and 70B versions of Llama 2 (touvron2023llama) along with Gemma3-4B (kamath2025gemma); in additional experiments we use Phi-2 (javaheripi2023phi) and Gemma3-1B. We use these models with either zero- or few-shot prompting, with both including appropriate task prompting such as “Classify the sentiment of the following sentence as positive or negative”. We denote our four Llama 2 model configurations as 7B-zero, 70B-zero, 7B-few and 70B-few, and our two Gemma 3 model configurations as Gemma3-4B-zero and Gemma3-4B-few.
(a) Llama 2 target models and Llama 2 surrogate models
(b) Llama 2 target models and Gemma3-4B surrogate model
Figure 2: Cheap surrogate models support effective data acquisition for active testing. We compare uniform-random testing to active testing (LURE) across four datasets. To guide active data acquisition we use a surrogate model that we train using a single step of in-context learning and then keep fixed. This stripped-back surrogate-model training, along with the use of small surrogate models relative to the target models, allows us to drastically reduce the computational cost of active testing while achieving strong performance.
We compare active testing to uniform-random testing, a standard approach to model evaluation. Acquisition methods from the active-learning literature (settles2012active) are not tailored for risk estimation (kossen2021active) and so would have limited utility as baseline methods here.
We use a logarithmic loss, as discussed in Sections 2 and 3. Like in past work (kossen2021active; kossen2022active), we measure the performance of a testing method using the median squared error of its risk estimate across random-number seeds, where the ground-truth risk is computed using a held-out test set. Alongside plots of the absolute value of this risk-estimation error, we present plots of relative error, namely the active risk-estimation error divided by the uniform-random risk-estimation error. Visualising performance on this relative scale lets us see the benefits of active testing more clearly. When we use a few-shot surrogate model with $n$ labelled in-context examples to evaluate a zero-shot target model, we ensure a fair comparison to uniform-random testing by comparing our results for $k$ acquired labels to uniform-random results for $n+k$ acquired labels.
5.2 Fixed surrogate models trained by in-context learning support effective data acquisition
First we investigate whether our switch from gradient-based training of the surrogate model to a single step of in-context learning still allows accurate risk estimation. We do this by running data acquisition across four datasets, four target models and three testing methods. In particular we compare uniform-random testing against two versions of active testing, LURE-7B and LURE-70B, differing only in the surrogate model used (7B-few vs 70B-few).
We find that active testing performs very well relative to the standard uniform-random testing in the vast majority of cases (Figure 2(a)), with risk-estimation error lowered by 32% on average (median over all the experimental variables listed above, as well as over possible label budgets between 1 and 400). This is an exciting result: a practically straightforward technique allows us to improve on standard practice for LLM evaluation with minimal computational overhead.
Contrasting with the overall positive result in Figure 2(a), LURE-70B performs poorly in evaluating the 70B-few model on the SST-2 dataset. We explore this failure case in Section 5.8 and show it is due to a large number of incorrect labels in the dataset. In Section C.5 we present an additional experiment that compares dynamic surrogate models to fixed ones: for the scenarios we study, dynamic surrogate models provide no accuracy benefits but incur prohibitive computational costs.
Figure 3: Older Llama 2 models are useful surrogates for active testing of newer Gemma 3 target models.
Figure 4: Our sampling-based active testing (LURE) approach works more reliably than interpolation-based active testing (ASE) when using cheaply constructed surrogate models. In ASE the surrogate model not only guides data acquisition but also much more directly affects the risk estimate: the labels used to compute the expected loss of the target model are simulated from the surrogate model.
5.3 Small surrogate models can be used to evaluate larger target models
A key additional result to highlight in Figure 2(a) is that LURE-7B works well not just for evaluating a target model of the same size, but also for evaluating a target model ten times larger. In fact, we show that we can reduce surrogate-model costs even further by using Gemma3-4B (Figure 2(b)) and Phi-2 (Figure 10(a)) to evaluate Llama 2 models, reducing computational costs by using even smaller surrogates. Meanwhile, older Llama 2 models effectively evaluate newer Gemma 3 models (Figure 3), which perform strongly on these datasets (Tables 3 and 2). As discussed in Section 3.2, this ability to use a small surrogate model—combined with the fact that fixing the surrogate model allows us to only compute one forward pass on the pool set—is a useful way of reducing costs.
5.4 Using stripped-back surrogate models favours sampling-based active testing
Next we evaluate our claim in Section 3.1 that sampling-based active testing is preferable over interpolation-based active testing given that the surrogate model is fixed. We do this by comparing our approach with ASE in evaluating the 7B-few and 70B-few models. Our results show ASE produces strong results on the HS dataset with the 70B-few model, but otherwise significantly underperforms (Figure 4). This is consistent with our explanation that ASE is more sensitive to the quality of the surrogate model, and that computational constraints on the surrogate model make it less practically useful. Note that the ASE risk estimator is constant if the surrogate model is fixed, as it is here.
5.5 Good data acquisition is possible without making predictions with the target model
Now we explore our suggestion in Section 3.3 that the surrogate model can be used to approximate the target model, such that we do not require target-model predictions during data acquisition. For the logarithmic loss we are using, this equates to using the predictive entropy of the surrogate model as the acquisition function. We assess this approach as applied to evaluating the 70B-few model with LURE-7B, and find it is surprisingly effective (Figure 6). This is promising: it suggests all the cost-reduction measures proposed in Section 3 are compatible with strong active-testing performance.
Figure 5: Using the surrogate model to approximate the target model during data acquisition (here this means acquiring data based on the predictive entropy of the surrogate model) reduces the need for target-model predictions while still performing well.
Figure 6: Active testing with a label-aware acquisition function (here the negative log likelihood of the surrogate model) enables dataset curation (selecting a subset of already-labelled test data) tailored to a target model of interest, helping reduce computational costs.
5.6 Active testing for dataset curation works, suggesting scope to reduce computational costs
Here we consider the alternative problem setting discussed in Section 3.4, where the inputs in the pool set are already labelled but it is too expensive to compute target-model predictions on all of them, with a possible solution being to curate a subset using a cheap surrogate model. More concretely we run active testing using the negative log likelihood as our label-aware acquisition function (Equation 4).
We find our approach improves over uniform-random subsampling on three datasets out of four (SST-2, Subj and HS) and performs slightly worse on FPB (Figure 6). These overall-positive results—along with those for the predictive-entropy function we considered in Section 5.5—suggest that active testing could be a useful tool for dataset curation, reducing computational costs when using existing evaluation datasets. The practical benefits of this could be significant: liang2023holistic required almost 20,000 hours of (Nvidia A100) GPU hours to evaluate 30 models on their HELM benchmark.
5.7 The benefit of active testing tends to hold in more challenging scenarios
Next we use the MMLU dataset to assess the robustness of active testing to an increase in task difficulty, which corresponds to worse surrogate-model performance (Tables 3 and 2) and thus poses a challenge for effective data acquisition. Our results show active testing continuing to outperform uniform-random testing for most combinations of surrogate and target models, with the only failure occurring when using 70B-few for both the surrogate and target models (Figure 7).
Figure 7: Active testing outperforms uniform-random testing in most cases on the harder MMLU dataset.
Figure 8: The failure of LURE-70B in efficiently evaluating the 70B-few model on the SST-2 dataset (left) is resolved by filtering incorrectly labelled datapoints from the pool set (middle, right). Underlying this effect is the presence of incorrect labels in the original dataset: we find that inputs with high loss are often mislabelled.
5.8 Incorrect labels can cause problems for active testing
Now we investigate the failure of LURE-70B to effectively evaluate the 70B-few model on the SST-2 dataset (Section 5.2). We construct two modified versions of the SST-2 pool set: one filtered to remove inputs for which 70B-few’s negative log likelihood (NLL) is greater than 5, and one filtered with a NLL threshold of 3. Then we re-run active testing with these modified pool sets.
We show that the failure case is reduced by the weaker filtering and completely resolved by the stronger filtering (Figure 8). To understand why, we inspect the filtered-out data and find it is often mislabelled. We find 75% of the 45 examples with NLL greater than 5 are mislabelled, and 67% of the 136 examples with NLL greater than 3. This is a useful demonstration of how active testing can fail when the source of labels is unreliable. Interestingly it can also be seen as a case for using less powerful surrogate models to guide data acquisition: LURE-7B worked well in the exact case where LURE-70B failed. It is also worth noting that in a dataset-curation setting (Section 5.6) we have access to all the test labels and therefore have the option to directly filter out high-NLL examples.
Figure 9: Our bootstrap estimator of the active-testing risk-estimation error provides a useful performance indicator. Coverage probability measures how often a confidence interval contains the true risk-estimation error.
5.9 Bootstrap estimation of risk-estimation error provides an accurate performance indicator
Having assessed our proposed active-testing method, we turn to our method for estimating the mean squared error of the LURE risk estimator (Section 4). We run active testing $T=100$ times, generating 100 sequences of acquired indices, $(i_{1:M})_{j=1}^{100}$. Then for $K\in\{1,2,\ldots,M\}$ we perform four steps:
  1. 1.
    Compute a ground-truth mean-squared error, $\mathrm{MSE}_{K}$, using Equation 5 with $q(i_{1:K})$ defined as a uniform distribution over the $K$ index sequences acquired so far.
  2. 2.
    Compute a bootstrap estimate of the MSE, $\widehat{\mathrm{MSE}}_{K}$, using Equation 6 with $B=1000$ estimates.
  3. 3.
    Compute the MSE-estimation error, $(\mathrm{MSE}_{K}-\widehat{\mathrm{MSE}}_{K})^{2}$.
  4. 4.
    Compute an approximate confidence interval, $\widehat{\mathrm{MSE}}_{K}\pm 2\hat{\sigma}$, using estimated standard deviation $\hat{\sigma}$.
Our results show MSE-estimation error converging to low values for most runs, after an initial period ($K<100$) of relatively high error (Figure 9). In addition to this, our confidence intervals provide high levels of coverage: for a strong majority of runs we see coverage probabilities of around 94% for $K\geq 100$, suggesting a practitioner can expect their computed confidence interval to contain the true mean squared error approximately 94% of the time. While there are cases where MSE estimation fails, the overall performance is promising, suggesting our proposed method could be a useful practical performance indicator, helping justify real-world deployment of active testing.
6 Related work
The sampling-based approach to active testing that our approach builds on was proposed by kossen2021active, using risk estimators from farquhar2021statistical. Earlier work on active testing proposed alternative methods. bennett2010online, ji2021active, katariya2012active and kumar2018classifier explored stratification-based techniques, such as splitting the pool set into strata (based on a measure of model confidence) and sampling uniformly within each stratum. sawade2010active and yilmaz2021sample proposed importance-sampling and Poisson-sampling methods related to that of kossen2021active but without the latter’s innovations of adaptive, surrogate-based data acquisition and risk estimation that accounts for sampling without replacement. nguyen2018active studied a special case of active testing focused on human vetting of noisy labels.
Recent years have seen various extensions of this earlier active-testing work. huang2025active introduced a clustering-based approach that relies on making predictions with the target model on the pool set, which we have argued is a barrier to scaling up. su2024metaat proposed a technique tailored to “dense” recognition tasks in computer vision (e.g., image segmentation and object detection). yu2024actively incorporated active testing into the process of training a model. ashurytahan2024label and hara2024active proposed methods for model selection.
The idea of reducing model-evaluation costs by using carefully selected test datasets has been highlighted a number of times in recent work. maynez2023benchmarking found that the datasets they studied could be substantially reduced in size, through uniform-random subsampling, while maintaining stable rankings of the models they were comparing. polo2024tinybenchmarks, saranathan2024dele and vivek2024anchor demonstrated more sophisticated methods for achieving the same goal. The goal in these studies was to take an existing test dataset (with labels for all inputs) and reduce it in a way that supports comparisons between different models. This prior work is therefore complementary to our contribution: we primarily study the problem of acquiring new data with which to evaluate a model of interest; and we use a method that tailors data acquisition to that given model.
7 Conclusion
We have argued that a significant barrier to principled active testing of large models is the computational cost of deciding which labels to acquire. To address this we have identified key contributors to the cost—training the surrogate model, and making predictions with the surrogate and target models—and have proposed straightforward cost-saving measures. We have shown that these measures are compatible with effective active testing, producing low-error risk estimates for large language models. On top of this we have demonstrated an estimator of risk-estimation accuracy that can provide an on-the-fly indication of how well active testing is working on a given practical problem. Overall we believe this represents substantial progress in more efficiently evaluating LLMs.
Acknowledgements
Freddie Bickford Smith was supported by the EPSRC CDT in Autonomous Intelligent Machines and Systems (EP/L015897/1). Tom Rainforth was supported by EPSRC grant EP/Y037200/1.
NeurIPS paper checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We support the claims in Sections 3, 4 and 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We highlight failure cases in Section 5, particularly in Section 5.8.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not present new theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide pseudocode in Algorithms 1 and 2, implementation details in Appendix B and code at github.com/gabrielleberrada/scaling-up-active-testing.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide pseudocode in Algorithms 1 and 2, implementation details in Appendix B and code at github.com/gabrielleberrada/scaling-up-active-testing. All datasets and models used in our experiments are open-source.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify data splits, hyperparameters and other implementation details in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report error bars and coverage probabilities in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide information about computational resources in Section B.7.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics and confirm that this research adheres to the code in all respects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss this in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We credit the creators of the datasets and models we use in Sections 5 and B.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: [N/A]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: [N/A]
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Impact statement
This paper presents work whose goal is to generally advance label- and compute-efficient model evaluation. As this is not limited to a particular area of application, it is difficult to evaluate the impact of this work. Generally speaking, we believe our work could play a role in reducing the computational costs, and thus the CO2 footprint, associated with model evaluation. As is often the case with research, we would caution against careless use of our method in real-world applications.
Appendix B Experiment details
B.1 Datasets
Our experiments are based on five datasets: Stanford Sentiment Treebank 2 (SST2; socher2013recursive; unknown license), Subjectivity (Subj; pang2004sentimental; Creative Commons Attribution 4.0 International License), Financial Phrase-bank (FPB; malo2013good; Creative Commons Attribution Non Commercial Share Alike 3.0 Unported License), Hate Speech (HS; degibert2018hate; Creative Commons Attribution-Share Alike 3.0 Spain License) and Massive Multitask Language Understanding (MMLU; hendrycks2021measuring; MIT license).
Each method, including uniform-random testing, acquires data from the same pool set (a set of candidate data for acquiring, with labels hidden until they are acquired), although its size may vary between experiments. We compare risk estimates against a ground-truth risk, which is computed on a separate test dataset in all cases except for FPB, where we use the pool set due to the small size of the original dataset. Table 1 summarizes sizes of the pool and test sets for each dataset.
SST-2 Subj FPB HS MMLU
Pool set 10,000 6,000 2,200 6,000 7,000
Test set 10,000 4,000 4,000 7,000
Table 1: Pool-set and test-set sizes for each dataset. Sets are randomly sampled so that they are disjoint.
Stanford Sentiment Treebank 2 (SST-2)
The Stanford Sentiment Treebank 2 dataset consists of 69,000 human-annotated sentences extracted from movie reviews. Each sentence is labeled as either “positive” or “negative”, providing a benchmark for assessing sentiment-analysis capabilities. We use the predefined split of the dataset, concentrating on the training set, which contains 56% positive and 44% negative labels. We randomly select two subsets of 10,000 sentences each from the training set to form the pool and test sets. The instruction used for this dataset is
[…]
The string "Answer" is replaced by "Label" for this dataset only.
Subjectivity (Subj)
The Subjectivity dataset is composed of 5,000 subjective sentences from movie reviews and 5,000 objective sentences from plot summaries. This dataset is used to determine sentiment polarity. The task is to classify each sentence as either “subjective” or “objective”. This dataset is balanced, with 49.5% of objective and 50.5% of subjective sentences. We randomly split the 10,000-sentence train set into a 6,000-sentence pool set and a 4,000-sentence test set. Examples for in-context-learning are selected from the separate 2,000-sentence test set. The instruction used for this dataset is
[…]
Financial Phrase-bank (FPB)
The Financial Phrase-bank dataset consists of 4,800 English financial news articles that were classified as “positive”, “neutral” and “negative” by human experts. We select the training subset of 2,200 sentences for which all 16 annotators agreed. This set contains 13% of negative, 25% of positive and 61% of neutral sentences. Due to its reduced size, the whole set is used as the pool set, and the approximation of the exact loss is the loss over the whole set. The instruction used for this dataset is
[…]
Hate Speech (HS)
The Hate Speech dataset is composed of 9,900 sentences extracted from Stormfront, a white-supremacist forum, from which we select the 9,600 sentences which are classified as “hate” or “no hate” to obtain a binary-classification task. The dataset is highly unbalanced, with 11% classified as hate speech and 89% as non-hate speech. We randomly split this dataset into a pool set of size $\sim$6,000 and a test set of size $\sim$4,000. The instruction used for this dataset is
[…]
Massive Multitask Language Understanding (MMLU)
The Massive Multitask Language Understanding dataset is composed of 116,000 sentences across 57 tasks asssing world knowledge and problem-solving. Each question is multiple-choice, with four choices available. We randomly split this dataset into a pool set of size 7,000 and a test set of size 7,000. The instruction used for this dataset is
[…]
B.2 Models
We use the 7B and 70B models from the Llama 2 family (touvron2023llama; Llama 2 Community License). We use 8 bit-quantisation for the 7B model and half-precision floating-point numbers for the 70B model; kossen2024context showed that these approximations do not significantly affect performance. We additionally use Gemma-3 4B (kamath2025gemma; Gemma License).
B.3 Prompt formatting
To evaluate a model on a classification task, we follow the formatting guidelines from kossen2024context in how we generate the input sentence. We begin with a dataset-specific instruction, such as
[…]
Then we introduce the sentence to be classified and ask for the corresponding label. The prompt is formatted as
[…]
B.4 In-context examples
We randomly select in-context examples such that each class is represented in proportion to its actual proportion in the dataset, which we found improved the model’s accuracy. These examples are included in the input between the instruction and the prompt. Each input example is formatted as
[…]
All few-shot models are given 50 in-context examples for all datasets, except for the MMLU dataset, for which models receive only 5 examples due to token limit. Examples are ordered randomly once and fixed for all evaluations, ensuring that all models receive the same context. The set of in-context examples is therefore fixed beforehand and is common to both the target model and the surrogate model, for label efficiency and fair comparison.
B.5 Generating and processing model outputs
To obtain deterministic, reproducible token generation, we set the maximum number of tokens to 1, do not set a top-$k$ nor top-$p$ value and output the logits directly, from which we compute probabilities.
For each input, the model outputs a logit value for each token in the vocabulary. We select the logits corresponding to the labels of the dataset and apply the softmax function to obtain a final probability distribution over the possible labels. Any logit that does not correspond to one of the labels is ignored.
In cases where a model represents a task label with multiple tokens (kossen2024context), we select only the relevant token that corresponds to the core of the word. For instance, the Llama 2 tokeniser encodes the word “objective” as [12091] but “subjective” as [4967, 573], which are tokens for “subject” and “ive”. In this case we select [12091] for “objective” and [4967] for “subjective”.
B.6 Data acquisition
Following kossen2021active, we clip values of the acquisition function to ensure no zero-mass inputs and thus guarantee that the LURE is unbiased. All acquisition probabilities below $\alpha=0.1$ times the probability corresponding to uniform-random acquisition are brought up to this limit value.
B.7 Computational resources
Our experiments are designed to be computationally efficient and therefore do not require substantial resources. Generating model outputs over the pool set is the main computational cost, typically requiring two GPUs for Llama 2 70B and one GPU for all other models used in this work. Aside from this, the active-testing procedure itself can be run efficiently on a small number of CPUs.
Appendix C Additional results
C.1 Model performance
Here we present accuracy and loss values for Llama2-7B, Llama2-70B and Gemma3-4B on the five datasets we study. Results for Llama 2 here match those published in kossen2024context.
Dataset L-7B-zero L-7B-few L-70B-zero L-70B-few G-4B-zero G-4B-few
FPB 66.75 90.60 26.25 94.15 29.70 91.00
SST-2 63.96 92.19 76.80 93.60 57.79 91.68
Subj 49.38 89.72 54.85 96.12 50.40 93.73
HS 11.38 89.28 84.18 90.23 42.87 89.42
MMLU 35.41 42.16 60.80 65.34 52.61 57.83
Table 2: Accuracy (%) of Llama 2 (L) 7B and 70B models and Gemma3-4B (G) model evaluated on pool sets.
Dataset L-7B-zero L-7B-few L-70B-zero L-70B-few G-4B-zero G-4B-few
FPB 0.7847 0.3084 1.3123 0.2152 1.4695 0.2701
SST-2 0.5639 0.2050 0.5808 0.1828 0.6234 0.2107
Subj 0.7326 0.3278 0.5873 0.1235 0.7101 0.2741
HS 0.9129 0.2979 0.5386 0.2166 0.7245 0.2678
MMLU 1.3468 1.2568 0.9217 0.8384 1.0920 0.9683
Table 3: Loss of Llama 2 (L) 7B and 70B models and Gemma3-4B (G) model evaluated on pool sets.
C.2 Failure mode on Subjectivity
Dataset 70B (T) & 70B (S) 70B (T) & 7B (S) 7B (T) & 70B (S) 7B (T) & 7B (S)
FPB 0.628 0.392 0.707 0.588
SST-2 0.285 0.224 0.527 0.454
Subj 0.364 – 0.025 0.802 0.652
HS 0.461 0.146 0.555 0.384
Table 4: Pearson’s correlation coefficient between the cross-entropy of the surrogate (S) and target (T) models’ predictions and the negative log likelihood of the target model’s predictions. Low or negative correlation indicates poor alignment between surrogate-based acquisition and optimal acquisition. Failure cases are in bold.
Here we build on our failure analysis in Section 5.8 (where we investigated why LURE-70B failed for the 70B-few target model on the SST-2 dataset), now focusing on the underperformance of LURE-7B for the 70B-few target model on the Subj dataset. We find that filtering out high-loss examples does not improve performance like it did in the SST-2 case, but we do find some insight from another line of investigation. In particular we compute the Pearson correlation coefficient between the cross-entropy-based acquisition scores and the optimal acquisition function (negative log likelihood of the target model’s predictions), with this correlation quantifying how well surrogate-based acquisition approximates optimal acquisition. Notably this is the only setting where the correlation is close to zero and is negative. This indicates that the surrogate model provides no meaningful guidance, helping to explain why active testing performs similarly to uniform-random testing in this case.
C.3 Smaller surrogate models on the four core datasets
(a) Llama 2 target models and Phi-2 surrogate model
(b) Llama 2 target models and Gemma3-1B surrogate model
Figure 10: Even very small surrogate models can support effective active testing in simple cases, although they struggle with evaluating our best-performing target model, 70B-few.
Here we explore two surrogate models that are smaller than those used in Section 5: Phi-2 2.7B (javaheripi2023phi; MIT License) and Gemma3-1B (kamath2025gemma; Gemma License). We see both surrogate models producing strong performance in evaluating 70B-zero and 7B-zero, and more mixed performance in evaluating the two stronger target models, 70B-few and 7B-few (Figure 10).
C.4 Smaller surrogate model on MMLU
Figure 11: A sufficiently weak surrogate model can cause active testing to consistently fail on MMLU.
Here we revisit Gemma3-1B as a surrogate model, now on the harder MMLU dataset. We see active testing failing to outperform uniform-random testing (Figure 11). This can be explained by how poorly Gemma3-1B performs on MMLU: it achieves approximately 25% accuracy, the base accuracy produced by uniform-random prediction. A weak surrogate model can thus undermine active testing.
C.5 Dynamic-surrogate active testing
Here we explore the use of dynamic surrogate models, in contrast with the fixed surrogate models we use in our approach. In the dynamic approach the surrogate model’s context incorporates newly acquired labels, leading to recomputed predictions, and thus novel acquisition probabilities at each sampling step. Our experiments focus on few-shot Llama2-7B and Llama2-70B target models with dynamic and fixed Llama2-7B surrogate models. We start with 10 in-context examples and run 40 steps of data acquisition. We run for 50 random-number seeds for dynamic-surrogate active testing, and for 3,000 seeds for fixed-surrogate active testing and uniform-random active testing.
Despite being much more computationally expensive, dynamic-surrogate active testing shows no consistent improvement in risk estimation over fixed-surrogate active testing (Tables 5 and 6). Thus, while dynamic surrogate models might be useful elsewhere, the results we see for our selection of datasets support our design choice of fixing the surrogate model: it improves over uniform-random testing while maintaining a low computational cost and thus readily scaling to LLM evaluations.
Dataset Testing method Step 10 Step 20 Step 30 Step 40
SST-2 Uniform random 91.41 45.49 29.25 20.66
(87.04, 95.86) (43.53, 47.52) (28.00, 30.51) (19.82, 21.52)
LURE-7B dynamic 63.25 22.79 13.52 8.62
(36.23, 96.67) (14.56, 32.91) (8.93, 18.80) (5.82, 11.88)
LURE-7B fixed 42.71 19.99 12.42 8.74
(40.32, 45.11) (19.07, 20.96) (11.87, 13.00) (8.38, 9.11)
FPB Uniform random 135.75 65.14 40.06 27.99
(130.00, 141.72) (62.37, 68.05) (38.41, 41.76) (26.84, 29.16)
LURE-7B dynamic 65.73 36.40 18.33 11.53
(45.79, 86.86) (25.31, 49.18) (11.82, 25.79) (7.37, 16.43)
LURE-7B fixed 77.54 36.88 22.97 16.26
(73.39, 81.94) (35.22, 38.61) (21.99, 23.95) (15.56, 16.96)
HS Uniform random 156.59 96.51 77.32 66.84
(152.02, 161.10) (93.78, 99.32) (75.14, 79.54) (64.93, 68.76)
LURE-7B dynamic 121.45 75.61 52.12 43.83
(100.99, 142.89) (60.04, 91.73) (40.48, 64.84) (35.35, 52.63)
LURE-7B fixed 144.44 89.65 72.62 63.67
(139.38, 149.87) (87.31, 91.98) (70.73, 74.47) (62.07, 65.31)
Subj Uniform random 46.97 21.79 13.76 10.23
(44.97, 49.00) (20.89, 22.70) (13.19, 14.36) (9.80, 10.66)
LURE-7B dynamic 71.76 35.25 36.75 30.94
(50.93, 95.54) (20.62, 53.32) (21.11, 55.19) (18.43, 46.09)
LURE-7B fixed 40.20 19.69 12.95 9.12
(38.39, 42.02) (18.89, 20.48) (12.42, 13.50) (8.73, 9.52)
Table 5: Mean squared error (multiplied by $10^{4}$; mean on first row; 5th and 95th percentile on second row) for uniform-random testing, dynamic-surrogate active testing and fixed-surrogate active testing of 7B-few.
Dataset Testing method Step 10 Step 20 Step 30 Step 40
SST-2 Uniform random 81.75 40.04 25.01 17.92
(77.77, 85.85) (38.24, 41.87) (23.98, 26.06) (17.19, 18.66)
LURE-7B dynamic 28.76 17.21 12.62 8.57
(22.13, 35.70) (11.42, 23.89) (8.83, 16.74) (6.37, 11.07)
LURE-7B fixed 40.09 19.11 12.26 8.44
(38.28, 41.98) (18.35, 19.89) (11.80, 12.73) (8.12, 8.76)
FPB Uniform random 173.90 80.36 50.14 35.07
(164.57, 183.50) (76.55, 84.22) (47.77, 52.54) (33.40, 36.83)
LURE-7B dynamic 114.11 34.89 19.71 11.10
(84.88, 144.11) (25.00, 46.22) (13.20, 27.05) (7.68, 14.95)
LURE-7B fixed 150.45 70.59 45.85 33.25
(140.12, 161.06) (66.56, 74.69) (43.59, 48.21) (31.64, 34.85)
HS Uniform random 98.22 48.73 33.04 25.27
(94.33, 102.28) (46.95, 50.54) (31.84, 34.29) (24.36, 26.20)
LURE-7B dynamic 88.73 41.15 23.93 16.18
(64.35, 116.85) (31.37, 51.26) (18.57, 29.54) (12.40, 20.03)
LURE-7B fixed 74.57 39.31 26.87 19.93
(71.50, 77.75) (37.78, 40.87) (25.85, 27.89) (19.18, 20.69)
Subj Uniform random 109.53 66.96 51.41 42.97
(105.47, 113.54) (64.75, 69.11) (49.77, 53.10) (41.59, 44.36)
LURE-7B dynamic 209.81 134.48 72.61 64.43
(161.39, 260.62) (99.07, 172.06) (54.29, 92.06) (48.04, 82.24)
LURE-7B fixed 134.87 77.31 57.14 47.24
(129.79, 140.22) (74.56, 80.06) (55.20, 59.07) (45.64, 48.81)
Table 6: Mean squared error (multiplied by $10^{4}$; mean on first row; 5th and 95th percentile on second row) for uniform-random testing, dynamic-surrogate active testing and fixed-surrogate active testing of 70B-few.
C.6 Risk-estimation values
In addition to the plots shown in Section 5, we present numerical values for the risk-estimation error of uniform-random testing and active testing (with the cross-entropy acquisition function) at acquisition steps 50, 100, 200, 300 and 400 in Tables 7, 8, 9, 10 and 11.
Target model Testing method Step 50 Step 100 Step 200 Step 300 Step 400
70B-zero Uniform random 24.3373 11.5663 5.1944 3.4854 2.4122
LURE-Llama2-7B 8.0511 3.6772 1.7589 1.0463
LURE-Llama2-70B 4.2225 2.0045 0.98390 0.64870
LURE-Gemma3-4B 7.1132 3.5176 1.6571 1.0540
7B-zero Uniform random 9.4687 4.8302 2.1823 1.4524 1.0075
LURE-Llama2-7B 2.5167 1.3094 0.59430 0.37380
LURE-Llama2-70B 1.9294 0.96850 0.42460 0.27060
LURE-Gemma3-4B 1.1421 0.5849 0.3681 0.2341
70B-few Uniform random 10.5749 5.16310 2.48680 1.51150 0.977800
LURE-Llama2-7B 6.3984 3.2058 1.3719 0.89110 0.63880
LURE-Llama2-70B 5.0757 2.5057 1.1921 0.76740 0.54300
LURE-Gemma3-4B 5.0390 2.6267 1.3102 0.8344 0.5825
7B-few Uniform random 14.3272 7.2276 3.4608 2.3147 1.5246
LURE-Llama2-7B 8.9252 4.5853 2.3314 1.5486 1.1345
LURE-Llama2-70B 4.8174 2.1675 1.1186 0.65290 0.45770
LURE-Gemma3-4B 6.5254 3.7684 1.9489 1.2371 0.8556
Table 7: Risk-estimation error (multiplied by $10^{4}$; minimum per target model in bold) for the FPB dataset.
Target model Testing method Step 50 Step 100 Step 200 Step 300 Step 400
70B-zero Uniform random 2.1154 1.0992 0.56060 0.34910 0.26150
LURE-Llama2-7B 0.5386 0.3089 0.1551 0.1023
LURE-Llama2-70B 0.4191 0.2308 0.1255 0.08190
LURE-Gemma3-4B 0.5283 0.2790 0.1542 0.1042
7B-zero Uniform random 12.4455 5.49 2.9545 1.9694 1.362
LURE-Llama2-7B 2.7160 1.4229 0.66190 0.48410
LURE-Llama2-70B 2.3740 1.2437 0.60510 0.42530
LURE-Gemma3-4B 3.5371 1.6843 0.8699 0.6062
70B-few Uniform random 36.6017 16.8825 9.4717 6.3474 4.3089
LURE-Llama2-7B 24.8052 14.7678 7.9162 5.2615 4.0596
LURE-Llama2-70B 40.8407 30.077 21.3892 17.1808 14.7939
LURE-Gemma3-4B 26.3377 16.8812 9.4499 7.1826 5.7175
7B-few Uniform random 19.5177 9.5602 4.9347 3.1286 2.2497
LURE-Llama2-7B 14.2747 8.3696 4.6265 3.2263 2.3764
LURE-Llama2-70B 9.7216 6.4914 3.6886 2.8971 2.1722
LURE-Gemma3-4B 13.3504 7.0426 3.9896 2.8028 1.9958
Table 8: Risk-estimation error (multiplied by $10^{4}$; minimum per target model in bold) for the SST-2 dataset.
Target model Testing method Step 50 Step 100 Step 200 Step 300 Step 400
70B-zero Uniform random 1.2104 0.64740 0.30890 0.21120 0.15430
LURE-Llama2-7B 0.4688 0.2311 0.1113 0.07570
LURE-Llama2-70B 0.139 0.0758 0.0373 0.0251
LURE-Gemma3-4B 0.3809 0.1897 0.0887 0.0638
7B-zero Uniform random 9.7212 5.1028 2.4378 1.5612 1.1429
LURE-Llama2-7B 3.4150 1.6835 0.87130 0.58730
LURE-Llama2-70B 1.6958 0.89970 0.44150 0.29280
LURE-Gemma3-4B 2.8896 1.4536 0.7144 0.4787
70B-few Uniform random 15.5609 6.6315 3.7465 2.5614 2.1347
LURE-Llama2-7B 16.4075 8.2611 4.4942 3.1802 2.1987
LURE-Llama2-70B 8.2473 5.0366 3.0052 2.1589 1.6293
LURE-Gemma3-4B 12.615 6.7789 3.3058 2.2544 1.7645
7B-few Uniform random 7.4461 3.9287 1.8132 1.1308 0.87000
LURE-Llama2-7B 5.2522 2.6268 1.2944 0.84500 0.63670
LURE-Llama2-70B 1.8539 1.0921 0.62250 0.42070 0.30200
LURE-Gemma3-4B 4.1246 2.1545 1.0256 0.6868 0.4941
Table 9: Risk-estimation error (multiplied by $10^{4}$; minimum per target model in bold) for the Subj dataset.
Target model Testing method Step 50 Step 100 Step 200 Step 300 Step 400
70B-zero Uniform random 2.3826 1.1493 6.3630 0.39690 0.30590
LURE-Llama2-7B 0.8497 0.4371 0.2247 0.1524
LURE-Llama2-70B 0.5703 0.2996 0.1418 0.09900
LURE-Gemma3-4B 0.7508 0.3700 0.1822 0.1330
7B-zero Uniform random 4.3455 2.3304 1.3545 0.96420 0.80470
LURE-Llama2-7B 3.8653 1.9395 1.1326 0.86310
LURE-Llama2-70B 3.0157 1.7693 0.94190 0.73000
LURE-Gemma3-4B 3.5508 1.6981 0.9918 0.7766
70B-few Uniform random 24.5497 11.8593 5.8947 4.263 2.897
LURE-Llama2-7B 26.4253 13.5746 6.9755 4.249 3.0014
LURE-Llama2-70B 19.8693 11.2704 6.1626 3.983 2.9785
LURE-Gemma3-4B 19.4658 10.4853 5.473 3.6247 2.767
7B-few Uniform random 23.7706 11.7404 6.6989 4.6215 3.4943
LURE-Llama2-7B 19.6371 9.638 5.8037 4.0226 3.1452
LURE-Llama2-70B 13.4897 7.5196 4.1388 2.7654 2.2691
LURE-Gemma3-4B 16.6135 8.7585 4.8131 3.4505 2.8044
Table 10: Risk-estimation error (multiplied by $10^{4}$; minimum per target model in bold) for the HS dataset.
Target model Testing method Step 50 Step 100 Step 200 Step 300 Step 400
70B-zero Uniform random 45.2305 22.0405 11.5532 7.2541 5.288
LURE-Llama2-7B 57.137 28.4008 13.3008 8.4398 6.7655
LURE-Llama2-70B 30.4079 17.3273 9.5205 6.6253 4.9939
LURE-Gemma3-4B 30.894 16.3988 9.0229 5.6838 4.1935
7B-zero Uniform random 8.1557 4.1386 2.1924 1.4752 1.1799
LURE-Llama2-7B 6.6835 3.2991 1.6763 1.2791 0.9767
LURE-Llama2-70B 4.3678 2.6164 1.3063 0.9088 0.7330
LURE-Gemma3-4B 4.9507 2.6543 1.4700 0.9569 0.7622
70B-few Uniform random 80.9096 41.8724 20.7615 12.5686 8.9944
LURE-Llama2-7B 94.483 51.3035 27.2732 17.5476 11.713
LURE-Llama2-70B 84.9646 45.9222 23.3469 15.7314 12.5486
LURE-Gemma3-4B 71.571 34.8396 19.3527 12.9035 9.2558
7B-few Uniform random 20.5838 10.142 4.8003 3.0952 2.4424
LURE-Llama2-7B 16.2533 8.686 4.5997 3.047 2.2239
LURE-Llama2-70B 7.242 4.3427 2.1991 1.4639 1.1166
LURE-Gemma3-4B 9.6443 4.9703 2.3827 1.694 1.2064
Table 11: Risk-estimation error (multiplied by $10^{4}$; minimum per target model in bold) for the MMLU dataset.